### v2

# SGS-SLAM: Semantic Gaussian Splatting for Neural Dense SLAM

Abstract
Semantic understanding plays a crucial role in Dense Simultaneous Localization and Mapping (SLAM). Recent advancements that integrate Gaussian Splatting into SLAM systems have demonstrated its effectiveness in generating high-quality renderings. Building on this progress, we propose SGS-SLAM which provides precise 3D semantic segmentation alongside high-fidelity reconstructions. Specifically, we propose to employ multi-channel optimization during the mapping process, integrating appearance, geometric, and semantic constraints with key-frame optimization to enhance reconstruction quality. Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, and semantic segmentation. It outperforms existing methods by a large margin meanwhile preserves real-time rendering ability.
Keywords SLAM  $\cdot$ 3D Reconstruction  $\cdot$ 3D Semantic Segmentation
Figure 1: The illustration of the proposed SGS-SLAM. It employs multi-channel 2D inputs encompassing appearance, geometry, and semantic information, leveraging Gaussian Splatting and differentiable rendering for joint parameter optimization. SGS-SLAM delivers accurate 3D semantic mapping along with high-fidelity map reconstructions.
**footnotetext: These authors contributed equally to this work
1 Introduction
Dense Visual Simultaneous Localization and Mapping (SLAM) is a crucial problem in the field of computer vision. It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses in a real-time manner. Traditional visual SLAM systems Davison et al. (2007); Newcombe et al. (2011); Salas-Moreno et al. (2013); Mur-Artal et al. (2015) stand out in sparse reconstruction using point clouds and voxels, but fall short in dense reconstruction. To extract dense geometric information for high-quality representation, learning-based SLAM methods Bloesch et al. (2018); Sucar et al. (2020) have gained wild attention. They demonstrate proficiency in generating decent 3D global maps meanwhile exhibiting robustness on noises and outliers. In addition, drawing inspiration from the advancements in the neural radiance field (NeRF) Mildenhall et al. (2021), NeRF-based SLAM approaches Sucar et al. (2021); Zhu et al. (2022); Kong et al. (2023); Zhang et al. (2023); Li et al. (2023a); Wang et al. (2023) have made further progress. They excel in producing accurate and high-fidelity global reconstruction by capturing dense photometric information through differentiable rendering.
However, NeRF-based SLAM methods employ multi-layer perceptrons (MLPs) as the implicit neural representation of scenes, which introduces several challenging limitations. Primarily, MLP models struggle with over-smoothing issues at the edge of objects, leading to a lack of fine-grained details in the map. This challenge also brings difficulties on disentangling the representation of objects, making it non-trivial to segment, edit, and manipulate objects within the scene. Moreover, when applied to larger scenes, MLP models are prone to catastrophic forgetting. This means that incorporating new scenes can adversely affect the precision of previously learned models, thereby reducing overall performance. Additionally, NeRF-based methods are computationally inefficient. Since the entire scene is modeled through one or several MLPs, it necessitates extensive model tuning for adding or updating scenes.
In this context, as opposite to NeRF-based neural representation, our exploration shifts towards the volumetric representation based on the 3D Gaussian Radiance Field Kerbl et al. (2023). This approach marks a significant shift and offers notable advantages in the scene representation.
Benefits from its rasterization of 3D primitives, Gaussian Splatting exhibits remarkably fast rendering speeds and allows direct gradient flow to each Gaussian’s parameters. This results in an almost linear projection between the dense photometric loss and parameters during optimization, unlike the hierarchical pixel-sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models. Moreover, the direct projection capability simplifies the addition of new parameters to the Gaussian field as separate channels, thereby enabling dynamic multi-channel rendering. Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed-reality. This integration allows real-time rendering of appearance, depth, and semantic color.
When compared with neural implicit semantic SLAM systems, such as DNS-SLAM Li et al. (2023a) and SNI-SLAM Zhu et al. (2023), our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy. Leveraging these benefits, our method enable precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering. Furthermore, the utilization of explicit spatial and semantic information for identifying scene content can be instrumental in optimizing camera tracking. Particularly, we incorporate a two-level adjustment based on geometric and semantic criteria for key-frames selection. This process relies on recognizing objects that have been previously seen in the trajectory. Extensive experiments are conducted on both synthetic and real-world scene benchmarks. These experiments compare our method against both implicit NeRF-based approaches Yang et al. (2022); Zhu et al. (2022); Wang et al. (2023); Johari et al. (2023), and novel 3D-Gaussian-based methods Keetha et al. (2023), evaluating performance in mapping, tracking, and semantic segmentation.
Overall, our work presents several key contributions, summarized as follows:
  • We introduce SGS-SLAM, the first semantic dense visual SLAM system grounded in 3D Gaussians. SGS-SLAM employs an explicit volumetric representation, enabling swift and real-time camera tracking and scene mapping. More importantly, it utilizes 2D semantic maps to learn 3D semantic representations expressed by Gaussians. Compared with previous NeRF-based methods which offers over-smooth object edges, SGS-SLAM provides high-fidelity reconstruction and optimal segmentation precision.
  • In SGS-SLAM, semantic maps provide additional supervision for optimizing parameters and selecting key frames. We employ a multi-channel parameter optimization strategy where appearance, geometric, and semantic signals collectively contribute to camera tracking and scene reconstruction. Furthermore, SGS-SLAM utilizes these diverse channels for the key-frame selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory. This approach results in efficient and high-quality map reconstruction based on the chosen key frames.
  • Utilizing semantic representation, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation. SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects in the map in real-time. This is achieved by grouping Gaussians by specifying the semantic labels of the objects.
2 Related Work
2.1 Semantic SLAM
Semantic information is of great importance for SLAM systems Mur-Artal et al. (2015); He et al. (2023); Whelan et al. (2015); Qin et al. (2018), which is a crucial requirement for applications in robotics and VR or AR fields. Real-time dense semantic SLAM systems Salas-Moreno et al. (2013); Bloesch et al. (2018); Rosinol et al. (2020) can integrate semantic information into 3D geometric representations. Traditional semantic SLAM systems rely on explicit 3D semantic expressions, such as voxel Hermans et al. (2014), point cloud Narita et al. (2019), and signed distance field Narita et al. (2019). These methods have limitations in terms of reconstruction speed, high-fidelity model acquisition, and memory usage. Moreover, traditional methods cannot reasonably infer unknown areas.
2.2 Neural Implicit SLAM
Methods based on NeRF McCormac et al. (2018), which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods Chung et al. (2023); Li et al. (2023b, 2024); Zhou et al. (2024). iMAP Sucar et al. (2021) uses a single MLP for scene representation, which shows limitation in large-scale scenes. NICE-SLAM Zhu et al. (2022) uses pre-trained multiple MLPs for hierarchical scene representation. Co-SLAM Wang et al. (2023) combines pixel set-based key-frame tracking with one-blob encoding. Go-SLAM Zhang et al. (2023) uses Droid-SLAM as front-end tracking system and multi-resolution hash encoding Müller et al. (2022) for mapping meanwhile integrated loop-closure detection and global optimization. However, these methods cannot utilize semantic information in maps. NIDS-SLAM Haghighi et al. (2023) leverages the mature front-end tracking of ORB-SLAM3 Campos et al. (2021) and Instant-NGP Müller et al. (2022) for mapping but does not optimize joint semantic features for 3D reconstruction. DNS-SLAM Li et al. (2023a) proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features. SNI-SLAM Zhu et al. (2023), a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
2.3 3D Gaussian Splatting-SLAM
The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting Kerbl et al. (2023) enable higher efficiency and accuracy on sparse pixel bases. However, existing 3DGS-based SLAM systems Yan et al. (2023); Keetha et al. (2023); Huang et al. (2023) lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes. We fuse semantic features into geometry and appearance and integrate 3D semantic features during the key-frame selection in the tracking process. This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
3 Method
SGS-SLAM is a Gaussian-based semantic visual SLAM systems. section 3.1 introduces its multi-channel Gaussian representation for joint parameter optimization. Like previous SLAM techniques, our method can be split into two processes: tracking and mapping. Tracking process estimates camera pose of each frame while keeping the scene parameters fixed. Mapping optimizes the scene representations based on the estimated camera pose. section 3.2 explains the breakdown steps in details. In addition, section 3.3 presents scene manipulation as a case study for downstream tasks. fig. 1 shows an overview of our system.
3.1 Multi-Channel Gaussian Representation
The scene is represented using a gaussian influence function $f(\cdot)$ on the map, For simplicity, these Gaussians are isotropic, as proposed in Keetha et al. (2023):
$f^{\rm 3D}(x)=\sigma\exp\left(-\frac{\|x-\mu\|^{2}}{2r^{2}}\right)$ (1)
Here, $\sigma\in[0,1]$ indicates opacity, $\mu\in\mathbb{R}^{3}$ represents the center position, and $r$ denotes the radius. Each Gaussian also carries RGB colors $c_{i}=[r_{i}~{}b_{i}~{}g_{i}]^{T}$.
In order to optimize the parameters of Gaussians to represent the scene, we need to render the Gaussians into 2D images in a differentiable manner. We use the render method from Luiten et al. (2024), providing extended functionality of rendering depth in colors. It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinate. The center of the Gaussian $\mu$, radius $r$, and depth $d$ (in camera coordinates) is splatted using the standard point rendering formula:
$\mu^{\rm 2D}=K\frac{E_{t}\mu}{d},\quad r^{\rm 2D}=\frac{lr}{d},\quad d=(E_{t}\mu)_{z}$ (2)
where $K$ is the camera intrinsic matrix, $E_{t}$ is the extrinsic matrix capturing the rotation and translation of the camera at frame $t$, $l$ is the focal length. The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula Max (1995):
$C_{\rm pix}=\sum_{i=1}^{n}c_{i}f_{i,\rm pix}^{\rm 2D}\prod_{j=1}^{i-1}(1-f_{j,\rm pix}^{\rm 2D})$ (3)
The pixel-level rendered color $C_{\rm pix}$ is the sum over the colors of each Gaussian $c_{i}$ and weighted by the influence function $f_{i,\rm pix}^{\rm 2D}$ (replace the 3D means and covariance matrices with the 2D splatted versions), multiplied by an occlusion term taking into account the effect of all Gaussians in front of the current Gaussian. Similarly, the depth can be rendered as:
$D_{\rm pix}=\sum_{i=1}^{n}d_{i}f_{i,\rm pix}^{\rm 2D}\prod_{j=1}^{i-1}(1-f_{j,\rm pix}^{\rm 2D})$ (4)
where $d_{i}$ denotes the depth of each Gaussian. By setting $d_{i}=1$, we are able to calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view. This aspect of visibility is essential for camera pose estimation, as it relies on the current reconstructed map. Additionally, it is also employed in map reconstruction, where new Gaussians are introduced in pixels lacking sufficient information.
While acquiring 3D semantic information is challenging and usually demands extensive manual labeling, 2D semantic label is a more accessible prior. In our approach, we leverage 2D semantic labels, which are often provided in datasets or can be easily obtained using state-of-the-art methods. We assign distinct channels to the parameters of Gaussians to denote their semantic labels and colors. During the rendering process, the 2D semantic map can be rendered from the reconstructed 3D scene as follows:
$S_{\rm pix}=\sum_{i=1}^{n}s_{i}f_{i,\rm pix}^{\rm 2D}\prod_{j=1}^{i-1}(1-f_{j,\rm pix}^{\rm 2D})$ (5)
where $s_{i}=[r_{i}~{}b_{i}~{}g_{i}]^{T}$ denotes the semantic color associated with the Gaussian. This semantic color is optimized jointly with the appearance color and depth during the mapping process.
3.2 Tracking and Mapping
3.2.1 Camera Pose Estimation
Given the first frame, the camera pose is set to identity and use as the reference coordinates for the following tracking and mapping procedure. While assessing the camera pose of an RGB-D view at a new timestep, the initial camera pose is determined by adding a displacement to the previous pose, assuming constant velocity, as $E_{t+1}=E_{t}+(E_{t}-E_{t-1})$. Following this, the current pose is iteratively refined by minimizing the tracking loss between the ground truth color ($C_{\rm pix}^{GT}$), depth images ($D_{\rm pix}^{GT}$), and semantic map ($S_{\rm pix}^{GT}$) and their differentiably rendered views:
$\mathcal{L}_{\rm tracking}=\sum_{\rm pix}(Sil_{\rm pix}>T_{\rm sil})(\lambda_{D}|D_{\rm pix}^{GT}-D_{\rm pix}|+\lambda_{C}|C_{\rm pix}^{GT}-C_{\rm pix}|+\lambda_{S}|S_{\rm pix}^{GT}-S_{\rm pix}|)$ (6)
Here, only those rendered pixels with a sufficiently large silhouette are factored into the loss calculation. The threshold $T_{\rm sil}$ is designed to make use of the map that has been previously optimized and has high certainty to be visible in the current camera view.
3.2.2 Key-frames Selection and Weighting
During the tracking phase of SLAM systems, key frames are identified and stored simultaneously. These key frames, providing different views of objects, are critical for mapping to refine 3D scene reconstruction. SGS-SLAM captures and stores key frames at constant time intervals. Subsequently, key frames associated with the current frame are chosen based on geometric and semantic constraints. Specifically, we randomly select pixels from the current frame and extract their corresponding Gaussians $G_{\rm sample}$ in the 3D scene. These Gaussians, $G_{\rm sample}$, are then projected onto the camera views of key frames as $G_{\rm proj}$. The $G_{\rm proj}$ are evaluated based on the geometric overlap ratio:
$\eta=\frac{1}{\sum G_{\rm proj}}\sum_{n=i}\{G_{i}|0\leq width(G_{i})\leq W,0\leq height(G_{i})\leq H\}$ (7)
It represents the proportion of Gaussians captured within the camera view of the key frames. $W$ and $H$ are the width and height of the camera view. The candidates with $\eta$ lower than certain threshold $T_{\rm geo}$ are removed. After the initial geometric-based selection, a second filter is conducted based on semantic criteria. We discard key frames whose semantic maps $S_{\rm pix}$ are identical to the current frame’s semantic map, as indicated by a high mIoU score. This threshold, denoted as $T_{\rm sem}$ intents to enhance map optimization from varying viewpoints, preferring views with low mIoU overlap. The remaining candidates are randomly sampled to serve as the selected key-frames associated with the current frame. In addition, we compute an uncertainty score for each key frame, defined as $\mathcal{U}(t)=e^{-\tau t}$, with $t$ representing the timestamp of the key frame and $\tau$ being a decay coefficient. This uncertainty score is used to weight the mapping loss $\mathcal{L}_{\rm mapping}$. The intuition behind this is that key frames with a later timestamp index carry a higher uncertainty in reconstruction due to the accumulation of camera tracking errors along the trajectory.
3.2.3 Map Reconstruction
The scene is modeled using Gaussians across three distinct channels: (1) their mean coordinates represent the geometric information of the scene, (2) their appearance colors depict the scene’s visual appearance, and (3) their semantic colors indicate the semantic labels of objects. These parameters across the channels are jointly optimized during the process of Gaussian densification and optimization, while the camera pose, ascertained from tracking, remains fixed.
Starting with the first frame, all pixels contribute to initializing the map. In the process of map reconstruction at a new timestep, new Gaussians are introduced to areas of the map that are either insufficiently dense or display new geometry in front of the previously estimated map. The addition of new Gaussians is regulated by applying a mask to the pixels where either (ii) the silhouette value $Sil_{\rm pix}$ falls below a certain threshold, signifying a high uncertainty in visibility, or (ii) the ground-truth depth is much smaller than the estimated depth, suggesting the presence of new geometric entities.
After densification, the parameters of the map are optimized by minimizing the mapping loss as:
$\mathcal{L}_{\rm mapping}=\mathcal{U}\sum_{\rm pix}\lambda_{D}|D_{\rm pix}^{GT}-D_{\rm pix}|+\lambda_{C}\mathcal{L}_{C}+\lambda_{S}\mathcal{L}_{S}$ (8)
where $\mathcal{L}_{C}$ and $\mathcal{L}_{S}$ are weighted SSIM loss Kerbl et al. (2023) with respect to appearance image and semantic image:
$\mathcal{L}(I_{\rm pix})=\sum_{\rm pix}\alpha|I_{\rm pix}^{GT}-I_{\rm pix}|+(1-\alpha)(1-ssim(I_{\rm pix}^{GT},I_{\rm pix}))$ (9)
$\lambda_{D}$, $\lambda_{C}$, $\lambda_{S}$, and $\alpha$ are predefined hyperparameters, and $\mathcal{U}$ is the uncertainty score defined in section 3.2.2.
Compared to existing NeRF-based approaches that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping. This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality. Compared to recent Gaussian-based methods, SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering. This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
3.3 Scene Manipulation
Given that the scene is represented explicitly by Gaussians, it becomes feasible to directly edit and manipulate a targeted group of Gaussians. In our case, Gaussians groups are identified based on their semantic labels. The mapping process generates these Gaussians, as defined in eq. 1, allowing for further manipulation in the following manner:
$f^{\rm 3D}_{\rm edit}(G,\tilde{y})=M(G,\tilde{y})\cdot\Phi_{T}(f^{\rm 3D}(G),\tilde{y})$ (10)
where the edited Gaussians, $f^{\rm 3D}_{\rm edit}$, are influenced by the visibility mask $M$, transition function $\Phi_{T}$, and the Gaussian’s semantic label $\tilde{y}$. The visibility mask $M$ determines if the Gaussians should be retained (1) or removed (0) based on $\tilde{y}$. The transition function $\Phi_{T}$ applies a transformation to the Gaussian’s coordinates on selected $\tilde{y}$, enabling spatial manipulation.
4 Experiment
4.1 Experimental Setup
4.1.1 Datasets
We evaluate our method on both synthetic and real-world datasets. To compare with other neural implicit SLAM methods, we evaluate on 8 synthetic scenes from Replica dataset Straub et al. (2019) and real-world scenes from ScanNet Dai et al. (2017a) and ScanNet++ Yeshwanth et al. (2023) datasets. The ground-truth camera pose and semantic map of Replica are offered from simulation, and ground-truth camera pose of ScanNet is generated by BundleFusion Dai et al. (2017b). The ground-truth semantic label is provided by the dataset.
4.1.2 Metrics
We use PSNR, Depth-L1 (on 2D depth map), SSIM, and LPIPS to evaluate the reconstruction quality. For the evaluation of camera pose, we adopt the average absolute trajectory error (ATE RMSE). For semantic segmentation, we calculate mIoU score.
4.1.3 Baselines
We compare the tracking and mapping with state-of-the-art methods iMap Sucar et al. (2021), Vox-Fusion Yang et al. (2022), NICE-SLAM Zhu et al. (2022), Co-SLAM Wang et al. (2023), ESLAM Johari et al. (2023), and SplaTAM Keetha et al. (2023). For semantic segmentation accuracy, we compare with NIDS-SLAM Haghighi et al. (2023), DNS-SLAM Li et al. (2023a), and SNI-SLAM Zhu et al. (2023).
4.2 Evaluation of Mapping and Localization
We show quantitative measures of reconstruction quality using the Replica dataset Straub et al. (2019) in table 1. Our method demonstrates state-of-the-art performance. When compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset Straub et al. (2019), with key details accentuated using colorful boxes. The results demonstrate that our method delivers more high-fidelity and robust reconstructions, particularly by capturing more detailed features of the objects in the scene.
In fig. 2, we present the reconstruction results of three chosen scenes, where regions of interest are accentuated with boxes in various colors. Our method exhibits high-fidelity reconstruction outcomes. Specifically, for small, intricately textured objects like a clock, socket, books on a tea table, and a lamp, our approach shows remarkable accuracy over NeRF-based methods. This is because Gaussians are capable of representing objects with complex textures and surfaces. Furthermore, NeRF-based methods often struggle with the over-smoothing issue, resulting in blurred edges on objects. In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes. Compared with SplaTAM, which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applies geometric constraints to preserve accurate shapes. This combination enables our method to achieve thorough modeling of both objects and their surrounding environment. The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
Table 1: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset Straub et al. (2019). Our method demonstrates SOTA performances on all three metrics.
Methods Metrics Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
[…]
Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset Straub et al. (2019). The values represent the average outcomes across eight scenes. The results of baselines are retrieved from Zhu et al. (2023). Our method remarkably outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics. Note that for parameter size, we only count the parameters of the model.
[…]
iMAP 4.645 3.118 4.153 9.92 2.23 1.82
[…]
table 2 displays the tracking evaluation results on the Replica dataset Straub et al. (2019). Our method excels in achieving the highest level of depth L1 loss (cm) and minimal ATE error, surpassing baseline methods by 70% in terms of depth loss and 34% in terms of ATE RMSE (cm). This exceptional performance can be attributed to our precise scene reconstruction, which provides finely-detailed rendering results. The high-quality rendering, in turn, contributes to accurate camera pose estimation based on the established map by preventing incorrect geometric reconstruction, which could otherwise result in inaccurate tracking outcomes. Additionally, utilizing features from different channels of Gaussians, such as geometry, appearance, and semantic information, provides multiple levels of supervision, resulting in a more robust and accurate tracking capability.
4.3 Evaluation of Semantic Segmentation
table 3 shows a quantitative evaluation of our method in comparison to other neural semantic SLAM approaches. It’s worth noting that we only show four scenes because previous NeRF-based semantic models only reported results on these scenes. In comparison to these previous methods, SGS-SLAM demonstrates state-of-the-art performance, outperforming the initial baseline by more than 10%. Substantial enhancement highlights the crucial advantage of explicit Gaussian representation over NeRF-based aprroaches. Gaussians can precisely isolate object boundaries, resulting in highly accurate 3D scene segmentation. In contrast, NeRF-based methods often struggle to recognize individual objects and typically require complex muti-level model designs and extensive feature fusion. Our approach offers an unparalleled ability to identify 3D objects in decomposed representations, which can serve as 3D priors for tracking and mapping in future time steps, and is well-suited for further downstream tasks.
Table 3: Quantitative comparison of our method against existing semantic NeRF-based SLAM methods on the Replica dataset Straub et al. (2019). The baselines are limited to four scenes as their results are reported only for these. For each scene, we compute the average mIoU score by comparing the rendered and the ground-truth 2D semantic image in the training view. Our method significantly outperforms the NeRF-based approaches, achieving mIoU scores over 90%.
Methods Avg. mIoU [%$\uparrow$ Room0 [%$\uparrow$ Room1 [%$\uparrow$ Room2 [%$\uparrow$ Office0 [%$\uparrow$
NIDS-SLAM 82.37 82.45 84.08 76.99 85.94
[…]
4.4 Evaluation of Key Frame Optimization
Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset Dai et al. (2017a). The rendered views display the reconstructed bike and guitar captured within the trajectory. Our method outperforms baselines by a large margin primarily due to the integration of key-frame optimization and semantic constraints. Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
In real-world datasets, tracking errors tend to accumulate along a trajectory, making pose estimations at later timestamps less reliable. Such inaccuracies can compromise the quality of map reconstructions, negatively impacting the accuracy of a previously well-established scene. A case in point is scene0000 from the ScanNet dataset Dai et al. (2017a), where objects such as bike and guitar are revisited at early and late stages in the trajectory. Key frames from later in the sequence, influenced by inaccurate camera poses, can disrupt the previously accurate reconstructions of these objects. fig. 3 illustrates the novel-view evaluation for scene0000. In comparison to SplaTAM, depicted in the upper row, our method delivers more accurate reconstruction outcomes. The bike, garbage bin, and guitar are accurately rendered, meanwhile details are preserved. Our method facilitates the selection of key frames base on geometric and semantic constraints, incorporating an uncertainty weighting during the optimization of selected key frames. This strategy demonstrate its effectiveness in map optimization from different views meanwhile preventing the unreliable key frame with high uncertainty to significantly altering the earlier accurately reconstructed map.
4.5 Scene Manipulation
The obtained semantic mask within the 3D scene has a range of applications for subsequent tasks. As an illustrative example, we demonstrate an straightforward but efficient Gaussian editing method defined by eq. 10, which is crucial for enabling scene manipulation for robotics or mixed reality applications.
Figure 4: The case study on scene manipulation in room0 of the Replica dataset Straub et al. (2019). We show the capabilities for object removal and transformation by specifying semantic labels. SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a table, as well as moving and rotating the table together with all objects on it.
Utilizing the decoupled scene representation, in contrast to NeRF-based approaches that demand fine-tuning of the entire network, SGS-SLAM has the ability to choose specific objects within the scene by referencing their semantic mask while keeping the remainder of the well-trained, irrelevant environment fixed. As shown in fig. 4, for object removal, we can directly erase the Gaussians associated with the editing target, such has removing the table while preserving all the items on it. In addition, we can group objects by selecting their semantic masks and apply translation and rotation, such as moving and rotating both the table and the above objects to a different place. This editing capability requires no training or fine-tuning, making it readily available for downstream applications.
4.6 Ablation Study
We perform the ablation of GS-SLAM on the scene0000 of the ScanNet dataset Dai et al. (2017a) to evaluate the effectiveness of multi-channel feature supervision, and the key-frame optimization strategies.
4.6.1 Effect of Multi-channel Optimization
table 4 shows the ablation study on multi-channel parameter optimization employed in SGS-SLAM. The results reveal that our optimization strategy can significantly improve the localization and mapping performance. Specifically, the system without appearance color, although performs poorly, still maintains its capability for tracking and depth estimation. This can be attributed to the gradient flow resulting from semantic supervision. Despite the Gaussian rasterizer used in our system not supporting backward flow through depth rendering, this underscores the value of semantic supervision within our framework. The absence of depth data leads to the poorest depth estimation, highlighting the importance of geometric supervision. Furthermore, the absence of an input semantic map disables 3D semantic segmentation and remarkably diminishes the performance of tracking and mapping. Additionally, the silhouette threshold, essential for assessing scene visibility, is crucial for the system stability. Without this threshold, the system shows a significant decline in the effectiveness of tracking and mapping.
Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset Dai et al. (2017a). The comparison involves settings where appearance, depth, and semantic supervision are removed.
Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
w/o color image 7.44 24.59 68.19
w/o depth map 47.66 40.47 15.14 54.52
w/o semantic map 9.15 13.81 17.52
w/o silhouette threshold 29.12 357.48 12.06 28.07
w/ multi-channel inputs 6.66 11.26 19.47 70.27
4.6.2 Effect of Key-frame Optimization
table 5 presents the results of key-frame selection ablation. Our two-level key-frame selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance. Additionally, without incorporating uncertainty weighting, the system demonstrates decrease in performance compared to its full implementation.
Table 5: Ablation study of key-frame optimization on the scene0000 of the ScanNet dataset Dai et al. (2017a). The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
w/o geometric threshold 6.66 15.55 19.21 68.93
w/o semantic threshold 8.44 12.89 17.84 69.85
w/o uncertainty weighting 6.87 11.43 18.72 70.12
w/ key-frame selection 6.18 11.26 19.47 70.27
5 Conclusion and Limitations
We presented SGS-SLAM, the first semantic dense visual SLAM system based on the 3D Gaussian representation. We propose to leverage multi-channel parameter optimization where appearance, geometric, and semantic constraints are combined to enforce high-accurate 3D semantic segmentation, and high-fidelity dense map reconstruction meanwhile effectively produce a robust camera pose estimation. SGS-SLAM takes the advantages of optimal key-frame optimization, resulting on reliable reconstruction quality. Extensive experiments show that our method provide state-of-the-art tracking and mapping results, meanwhile maintain rapid rendering speeds. Furthermore, the high-quality reconstruction of scenes and precise 3D semantic labeling generated by our system establish a strong foundation for the downstream tasks such as scene editing, offers solid prior for the robotics or AR/VR applications.
5.1 Limitations
SGS-SLAM replies on depth and 2D semantic signal inputs for tracking and mapping. In scenarios where this information is scarce or difficult to access, the system’s effectiveness will be compromised. Additionally, our method incurs large memory consumption when deployed in large scenes. Addressing these limitations will be an objective for future research.
Appendix A Experiment Settings
In this section, we outline the experimental setup and hyperparameters applied in our studies.
A.1 SGS-SLAM Settings
By default, both mapping and tracking operations are conducted for each frame. During the tracking phase, we set the silhouette visibility threshold, $T_{\rm sil}$, to 0.99. The multi-channel optimization involves three parameters: $\lambda_{D}=1.0$ for depth, $\lambda_{C}=0.5$ for colors, and $\lambda_{S}=0.05$ for semantic loss, with the semantic loss weight being comparatively low due to the typical noisiness of real-world semantic labels. Throughout the tracking, the multi-channel Gaussian parameters remain constant, adjusting only the camera parameters with a learning rate of 2e-3 for transition. Key-frames are initially chosen at intervals of every 5 frames, then refined based on geometric and semantic criteria. The geometric overlap threshold, $\eta$, is defined at 0.05, and the semantic mean Intersection over Union (mIoU) threshold, $T_{\rm sem}$, at 0.7. The maximum number of key-frames per frame is limited to 25, considering the computation speed. The uncertainty decay coefficient, $\tau$ scales with the length of the input frame series. In mapping process, the silhouette threshold $T_{\rm sil}$ is adjusted to 0.5. The weights of photometric loss are set to $\lambda_{D}=1.0$, $\lambda_{C}=0.5$, and $\lambda_{S}=0.1$. Here, camera parameters are fixed, and Gaussian parameters are optimized, with specific learning rates for 3D position at 1e-4, color 2.5e-3, Gaussian rotation at 1e-3, logit opacity at 0.05, and logg scale at 1e-3. Performance metrics of tracking and mapping are assessed every 5 frames, with mIoU scores evaluated at the same frequency.
The mapping and tracking iteration steps are specific to each dataset, In the case of the Replica dataset Straub et al. [2019], the number of iterations for tracking and mapping are set to 40 and 60. For the ScanNet dataset Dai et al. [2017a], tracking and mapping are set to 120 and 40. In the enhanced ScanNet++ dataset Yeshwanth et al. [2023], where the camera transition is large between each frame, the tracking and mapping iterations are adjusted to 220 and 50.
A.2 Baseline Settings
We adhere to the default configurations for each baseline as reported in their papers. The evaluation metrics for tracking and mapping are consistent with those applied to our method. For baselines whose implementations are not publicly available, we present the results as reported in their papers.
Appendix B Additional Experiment Results
We provide additional quantitative analysis of camera tracking in section B.1. The visualization of semantic segmentation compared with NeRF-based method is presented in section B.2. More qualitative novel view rendering results are illustrated in section B.3. We compared our method with Vox-Fusion Yang et al. [2022], NICE-SLAM Zhu et al. [2022], Co-SLAM Wang et al. [2023], ESLAM Johari et al. [2023], and Point-SLAM Sandström et al. [2023] for ATE RMSE evaluation. For 3D semantic segmentation, we visualized the comparison with DNS-SLAM Li et al. [2023a].
B.1 Camera Tracking
In this section, we break down the quantitative analysis on ATE RMSE [cm] on Replica Straub et al. [2019], ScanNet Dai et al. [2017a], and ScanNet++ Yeshwanth et al. [2023] datasets. table 6, table 7, and table 8 present the evaluation our SGS-SLAM against baseline models on each dataset. Our method of estimating camera poses by directly optimizing the gradient on dense photometric loss achieves state-of-the-art tracking performance on datasets with high-quality RGB-D images. In particular, on the ScanNet++ dataset Yeshwanth et al. [2023], where there is a large camera transition between successive frames, NeRF-based methods like ESLAM failed to track. Conversely, SGS-SLAM demonstrated robust tracking performance.
Table 6: Quantitative comparison of ATE RMSE [cm] between our method and the baselines for each scene of the Replica dataset Straub et al. [2019].
Methods Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
Vox-Fusion 3.09 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94
[…]
Table 7: Quantitative comparison of ATE RMSE [cm] between our method and the baselines for the selected scenes on the ScanNet dataset Dai et al. [2017a].
Methods Avg. $0$ 0059 0106 0169 0181 0207
Vox-Fusion 26.90 68.84 24.18 8.41 27.28 23.30 9.41
[…]
Table 8: Quantitative comparison of ATE RMSE [cm] between our method and the baseline for the selected scenes on the ScanNet++ dataset Yeshwanth et al. [2023].
Methods Avg. [cm]$\downarrow$ 8b5caf3398 [cm]$\downarrow$ b20a261fdf [cm]$\downarrow$
ESLAM 170.06 185.15 156.96
[…]
B.2 Semantic Segmentation
In this section, the outcomes of semantic segmentation on the Replica dataset Straub et al. [2019] are visualized and compared with DNS-SLAM Li et al. [2023a], a NeRF-based approach. As illustrated, our method offered accurate and detailed segmentation, whereas DNS-SLAM faces challenges in edges due to the over-smoothing issue of NeRF.
Figure 5: Qualitative comparison of our method and DNS-SLAM Li et al. [2023a] for semantic segmentation from the Replica dataset Straub et al. [2019]. The visualization outcomes of DNS-SLAM Li et al. [2023a] are obtained from its paper. The frames of training view are chosen based on the results presented in DNS-SLAM. Compared to NeRF-based models, our approach delivers segmentation results with higher accuracy.
B.3 Novel View Rendering
We present additional results of novel view rendering using our method across the Replica Straub et al. [2019], ScanNet Dai et al. [2017a], and ScanNet++ Yeshwanth et al. [2023] datasets, with comparisons to ESLAM Johari et al. [2023]. Visualizations are provided in fig. 6, fig. 7, fig. 8, and fig. 9 with semantic segmentation outcomes. Our method consistently delivers high-quality rendering results for both synthesized and real-world datasets. Notably, on the challenging real-world ScanNet++ dataset, ESLAM Johari et al. [2023] struggled to reconstruct the scene. By contrast, SGS-SLAM provides accurate high-fidelity scene reconstructions along with precise segmentation outcomes. Note that the ground-truth segmentation labels are retrieved from the ground-truth mesh at the instance level, and therefore, our results also show instance-level segmentation.
Figure 6: The visualization of novel view rendering between the ESLAM Johari et al. [2023] and our method on the Replica dataset Straub et al. [2019]. The ground-truth novel views are captured from meshes provided by the dataset.
Figure 7: The visualization of novel view rendering between the baseline and our method using the ScanNet dataset Dai et al. [2017a]. The ground-truth novel views are captured from meshes. SGS-SLAM exhibits rendering of high fidelity, outperforms the NeRF-based ESLAM Johari et al. [2023]. In contrast to the ground-truth mesh, our method demonstrates robust mapping in areas where the ground-truth mesh presents holes.
Figure 8: The visualization of 3D semantic segmentation results of SGS-SLAM, as applied to the novel views selected in fig. 7. Note that the rendering results exhibit minor variations in scene objects due to the use of an modified semantic dataset from ScantNet. For our method, the training data is processed from the filtered semantic labels using the nyu40-class, where certain objects are not distinctly labeled and are assigned as background (depicted in black). Furthermore, we introduce extra labels, like guitar, bag, and basket, to enhance the quality of scene reconstruction.
Figure 9: The visualization of novel view rendering between the baseline and our method using the ScanNet++ dataset Yeshwanth et al. [2023]. The ground-truth novel views are captured from meshes. SGS-SLAM demonstrates superior rendering quality, while ESLAM Johari et al. [2023] suffers from significant tracking errors and fails to reconstruct the map. In addition, our method also offers accurate instance-level segmentation outcomes.

### v3

@@
 Abstract
-Semantic understanding plays a crucial role in Dense Simultaneous Localization and Mapping (SLAM).
-Recent advancements that integrate Gaussian Splatting into SLAM systems have demonstrated its effectiveness in generating high-quality renderings.
-Building on this progress, we propose SGS-SLAM which provides precise 3D semantic segmentation alongside high-fidelity reconstructions.
-Specifically, we propose to employ multi-channel optimization during the mapping process, integrating appearance, geometric, and semantic constraints with key-frame optimization to enhance reconstruction quality.
-Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, and semantic segmentation.
-It outperforms existing methods by a large margin meanwhile preserves real-time rendering ability.
+We present SGS-SLAM, the first semantic visual SLAM system based on Gaussian Splatting.
+It incorporates appearance, geometry, and semantic features through multi-channel optimization, addressing the oversmoothing limitations of neural implicit SLAM systems in high-quality rendering, scene understanding, and object-level geometry.
+We introduce a unique semantic feature loss that effectively compensates for the shortcomings of traditional depth and color losses in object optimization.
+Through a semantic-guided keyframe selection strategy, we prevent erroneous reconstructions caused by cumulative errors.
+Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, precise semantic segmentation, and object-level geometric accuracy, while ensuring real-time rendering capabilities.
 Keywords SLAM  $\cdot$ 3D Reconstruction  $\cdot$ 3D Semantic Segmentation
 Figure 1: The illustration of the proposed SGS-SLAM.
-It employs multi-channel 2D inputs encompassing appearance, geometry, and semantic information, leveraging Gaussian Splatting and differentiable rendering for joint parameter optimization.
-SGS-SLAM delivers accurate 3D semantic mapping along with high-fidelity map reconstructions.
-**footnotetext: These authors contributed equally to this work
+It employs 2D inputs encompassing appearance, geometry, and semantic information, leveraging Gaussian Splatting and differentiable rendering for multi-channel parameter optimization.
+During the mapping process, SGS-SLAM maps the 2D semantic prior to the 3D scene, jointly optimizing it via the mapping loss for accurate 3D segmentation outcomes.
+**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author
 1 Introduction
@@
 Primarily, MLP models struggle with over-smoothing issues at the edge of objects, leading to a lack of fine-grained details in the map.
-This challenge also brings difficulties on disentangling the representation of objects, making it non-trivial to segment, edit, and manipulate objects within the scene.
+This challenge also brings difficulties in disentangling the representation of objects, making it non-trivial to segment, edit, and manipulate objects within the scene.
 Moreover, when applied to larger scenes, MLP models are prone to catastrophic forgetting.
@@
 Benefits from its rasterization of 3D primitives, Gaussian Splatting exhibits remarkably fast rendering speeds and allows direct gradient flow to each Gaussian’s parameters.
-This results in an almost linear projection between the dense photometric loss and parameters during optimization, unlike the hierarchical pixel-sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models.
+This results in an almost linear projection between the dense photometric loss and parameters during optimization, unlike the hierarchical pixel sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models.
 Moreover, the direct projection capability simplifies the addition of new parameters to the Gaussian field as separate channels, thereby enabling dynamic multi-channel rendering.
-Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed-reality.
+Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed reality.
 This integration allows real-time rendering of appearance, depth, and semantic color.
@@
 (2023), our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy.
-Leveraging these benefits, our method enable precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering.
-Furthermore, the utilization of explicit spatial and semantic information for identifying scene content can be instrumental in optimizing camera tracking.
-Particularly, we incorporate a two-level adjustment based on geometric and semantic criteria for key-frames selection.
+Leveraging these benefits, our method enables precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering.
+Furthermore, using explicit spatial and semantic information to identify scene content can be instrumental in optimizing camera tracking.
+Particularly, we incorporate a two-level adjustment based on geometric and semantic criteria for keyframes selection.
 This process relies on recognizing objects that have been previously seen in the trajectory.
@@
 More importantly, it utilizes 2D semantic maps to learn 3D semantic representations expressed by Gaussians.
-Compared with previous NeRF-based methods which offers over-smooth object edges, SGS-SLAM provides high-fidelity reconstruction and optimal segmentation precision.
+Compared with previous NeRF-based methods which offer over-smooth object edges, SGS-SLAM provides high-fidelity reconstruction and optimal segmentation precision.
   • In SGS-SLAM, semantic maps provide additional supervision for optimizing parameters and selecting keyframes.
 We employ a multi-channel parameter optimization strategy where appearance, geometric, and semantic signals collectively contribute to camera tracking and scene reconstruction.
-Furthermore, SGS-SLAM utilizes these diverse channels for the key-frame selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory.
+Furthermore, SGS-SLAM utilizes these diverse channels for keyframe selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory.
 This approach results in efficient and high-quality map reconstruction based on the chosen keyframes.
   • Utilizing semantic representation, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation.
-SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects in the map in real-time.
+SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects in the map in real time.
 This is achieved by grouping Gaussians by specifying the semantic labels of the objects.
@@
 (2018); Rosinol et al.
-(2020) can integrate semantic information into 3D geometric representations.
-Traditional semantic SLAM systems rely on explicit 3D semantic expressions, such as voxel Hermans et al.
+(2020) integrate semantic information into 3D geometric representations.
+Traditional semantic SLAM systems rely on sparse 3D semantic expressions, such as voxel Hermans et al.
 (2014), point cloud Narita et al.
@@
 (2019).
-These methods have limitations in terms of reconstruction speed, high-fidelity model acquisition, and memory usage.
-Moreover, traditional methods cannot reasonably infer unknown areas.
+These methods struggle with accurately interpreting complex environments due to limited semantic understanding.
+This results in a simplified categorization of environmental features, which may not capture the full range of objects and their relationships within a space.
+Moreover, these methods exhibit limitations regarding reconstruction speed, high-fidelity model acquisition, and memory usage.
 2.2 Neural Implicit SLAM
@@
 iMAP Sucar et al.
-(2021) uses a single MLP for scene representation, which shows limitation in large-scale scenes.
+(2021) uses a single MLP for scene representation, which shows limitations in large-scale scenes.
 NICE-SLAM Zhu et al.
@@
 Co-SLAM Wang et al.
-(2023) combines pixel set-based key-frame tracking with one-blob encoding.
+(2023) combines pixel set-based keyframe tracking with one-blob encoding.
 Go-SLAM Zhang et al.
-(2023) uses Droid-SLAM as front-end tracking system and multi-resolution hash encoding Müller et al.
+(2023) uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding Müller et al.
 (2022) for mapping meanwhile integrated loop-closure detection and global optimization.
@@
 (2023), a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
-2.3 3D Gaussian Splatting-SLAM
+2.3 3D Gaussian Splatting SLAM
 The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting Kerbl et al.
-(2023) enable higher efficiency and accuracy on sparse pixel bases.
+(2023) enable higher efficiency and accuracy on scene reconstruction.
 However, existing 3DGS-based SLAM systems Yan et al.
@@
 (2023) lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes.
-We fuse semantic features into geometry and appearance and integrate 3D semantic features during the key-frame selection in the tracking process.
+We fuse semantic features into geometry and appearance and integrate 3D semantic features during keyframe selection in the tracking process.
 This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
 3 Method
-SGS-SLAM is a Gaussian-based semantic visual SLAM systems.
+SGS-SLAM is a Gaussian-based semantic visual SLAM system.
 section 3.1 introduces its multi-channel Gaussian representation for joint parameter optimization.
 Like previous SLAM techniques, our method can be split into two processes: tracking and mapping.
-Tracking process estimates camera pose of each frame while keeping the scene parameters fixed.
+The tracking process estimates the camera pose of each frame while keeping the scene parameters fixed.
 Mapping optimizes the scene representations based on the estimated camera pose.
-section 3.2 explains the breakdown steps in details.
+section 3.2 explains the breakdown steps in detail.
 In addition, section 3.3 presents scene manipulation as a case study for downstream tasks.
@@
 (2024), providing extended functionality of rendering depth in colors.
-It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinate.
+It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinates.
 The center of the Gaussian $\mu$, radius $r$, and depth $d$ (in camera coordinates) is splatted using the standard point rendering formula:
@@
 where $d_{i}$ denotes the depth of each Gaussian.
-By setting $d_{i}=1$, we are able to calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view.
+By setting $d_{i}=1$, we can calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view.
 This aspect of visibility is essential for camera pose estimation, as it relies on the current reconstructed map.
 Additionally, it is also employed in map reconstruction, where new Gaussians are introduced in pixels lacking sufficient information.
-While acquiring 3D semantic information is challenging and usually demands extensive manual labeling, 2D semantic label is a more accessible prior.
-In our approach, we leverage 2D semantic labels, which are often provided in datasets or can be easily obtained using state-of-the-art methods.
+While acquiring 3D semantic information is challenging and usually demands extensive manual labeling, the 2D semantic label is more accessible prior.
+In our approach, we leverage 2D semantic labels, which are often provided in datasets or can be easily obtained using off-the-shelf methods.
 We assign distinct channels to the parameters of Gaussians to denote their semantic labels and colors.
@@
 This semantic color is optimized jointly with the appearance color and depth during the mapping process.
+The Gaussian representations employed in SGS-SLAM facilitate high-quality reconstructions at high rendering speed, offering exceptional accuracy in capturing complex textures and geometry with remarkable detail and efficiency.
+Furthermore, the integration of semantic features within our method significantly advances optimal scene interpretation and precise object-level geometry, effectively mitigating the oversmoothing issues prevalent in NeRF models.
 3.2 Tracking and Mapping
 3.2.1 Camera Pose Estimation
-Given the first frame, the camera pose is set to identity and use as the reference coordinates for the following tracking and mapping procedure.
+Given the first frame, the camera pose is set to identity and used as the reference coordinates for the following tracking and mapping procedure.
 While assessing the camera pose of an RGB-D view at a new timestep, the initial camera pose is determined by adding a displacement to the previous pose, assuming constant velocity, as $E_{t+1}=E_{t}+(E_{t}-E_{t-1})$.
@@
 The threshold $T_{\rm sil}$ is designed to make use of the map that has been previously optimized and has high certainty to be visible in the current camera view.
-3.2.2 Key-frames Selection and Weighting
+3.2.2 Keyframes Selection and Weighting
 During the tracking phase of SLAM systems, key frames are identified and stored simultaneously.
@@
 $W$ and $H$ are the width and height of the camera view.
-The candidates with $\eta$ lower than certain threshold $T_{\rm geo}$ are removed.
+The candidates with $\eta$ lower than a certain threshold $T_{\rm geo}$ are removed.
 After the initial geometric-based selection, a second filter is conducted based on semantic criteria.
 We discard keyframes whose semantic maps $S_{\rm pix}$ are identical to the current frame’s semantic map, as indicated by a high mIoU score.
-This threshold, denoted as $T_{\rm sem}$ intents to enhance map optimization from varying viewpoints, preferring views with low mIoU overlap.
-The remaining candidates are randomly sampled to serve as the selected key-frames associated with the current frame.
+This threshold, denoted as $T_{\rm sem}$ intends to enhance map optimization from varying viewpoints, preferring views with low mIoU overlap.
+The remaining candidates are randomly sampled to serve as the selected keyframes associated with the current frame.
 In addition, we compute an uncertainty score for each keyframe, defined as $\mathcal{U}(t)=e^{-\tau t}$, with $t$ representing the timestamp of the keyframe and $\tau$ being a decay coefficient.
@@
 The scene is modeled using Gaussians across three distinct channels: (1) their mean coordinates represent the geometric information of the scene, (2) their appearance colors depict the scene’s visual appearance, and (3) their semantic colors indicate the semantic labels of objects.
-These parameters across the channels are jointly optimized during the process of Gaussian densification and optimization, while the camera pose, ascertained from tracking, remains fixed.
+These parameters across the channels are jointly optimized during the process of Gaussian densification and optimization, whereas the camera pose, ascertained from tracking, remains fixed.
 Starting with the first frame, all pixels contribute to initializing the map.
@@
 The addition of new Gaussians is regulated by applying a mask to the pixels where either (ii) the silhouette value $Sil_{\rm pix}$ falls below a certain threshold, signifying a high uncertainty in visibility, or (ii) the ground-truth depth is much smaller than the estimated depth, suggesting the presence of new geometric entities.
-After densification, the parameters of the map are optimized by minimizing the mapping loss as:
+After densification, the parameters of the map are optimized by minimizing the mapping loss:
 $\mathcal{L}_{\rm mapping}=\mathcal{U}\sum_{\rm pix}\lambda_{D}|D_{\rm pix}^{GT}-D_{\rm pix}|+\lambda_{C}\mathcal{L}_{C}+\lambda_{S}\mathcal{L}_{S}$ (8)
@@
 $\lambda_{D}$, $\lambda_{C}$, $\lambda_{S}$, and $\alpha$ are predefined hyperparameters, and $\mathcal{U}$ is the uncertainty score defined in section 3.2.2.
-Compared to existing NeRF-based approaches that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping.
+Compared to existing NeRF-based approaches Zhu et al.
+(2022); Johari et al.
+(2023); Li et al.
+(2023a); Zhu et al.
+(2023) that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping.
 This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality.
-Compared to recent Gaussian-based methods, SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering.
+Compared to recent Gaussian-based methods Keetha et al.
+(2023); Yan et al.
+(2023), SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering.
 This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
-3.3 Scene Manipulation
+3.3 Scene Manipulation via Object-level Geometry
 Given that the scene is represented explicitly by Gaussians, it becomes feasible to directly edit and manipulate a targeted group of Gaussians.
-In our case, Gaussians groups are identified based on their semantic labels.
+In our case, Gaussian groups are identified based on their semantic labels.
 The mapping process generates these Gaussians, as defined in eq.
@@
 We evaluate our method on both synthetic and real-world datasets.
-To compare with other neural implicit SLAM methods, we evaluate on 8 synthetic scenes from Replica dataset Straub et al.
+To compare with other neural implicit SLAM methods, we evaluate synthetic scenes from Replica dataset Straub et al.
 (2019) and real-world scenes from ScanNet Dai et al.
@@
 (2023) datasets.
-The ground-truth camera pose and semantic map of Replica are offered from simulation, and ground-truth camera pose of ScanNet is generated by BundleFusion Dai et al.
+The ground-truth camera pose and semantic map of Replica are offered from simulation, and the ground-truth camera pose of ScanNet is generated by BundleFusion Dai et al.
 (2017b).
-The ground-truth semantic label is provided by the dataset.
+The ground-truth 2D semantic label is provided by the dataset.
 4.1.2 Metrics
-We use PSNR, Depth-L1 (on 2D depth map), SSIM, and LPIPS to evaluate the reconstruction quality.
+We use PSNR, Depth-L1 (on 2D depth maps), SSIM, and LPIPS to evaluate the reconstruction quality.
 For the evaluation of camera pose, we adopt the average absolute trajectory error (ATE RMSE).
@@
 We compare the tracking and mapping with state-of-the-art methods iMap Sucar et al.
-(2021), Vox-Fusion Yang et al.
 (2021), NICE-SLAM Zhu et al.
@@
 Our method demonstrates state-of-the-art performance.
-When compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
+Compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
 Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset Straub et al.
@@
 In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes.
-Compared with SplaTAM, which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applies geometric constraints to preserve accurate shapes.
+Compared with SplaTAM, which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes.
 This combination enables our method to achieve thorough modeling of both objects and their surrounding environment.
@@
 In comparison to these previous methods, SGS-SLAM demonstrates state-of-the-art performance, outperforming the initial baseline by more than 10%.
-Substantial enhancement highlights the crucial advantage of explicit Gaussian representation over NeRF-based aprroaches.
+Substantial enhancement highlights the crucial advantage of explicit Gaussian representation over NeRF-based approaches.
 Gaussians can precisely isolate object boundaries, resulting in highly accurate 3D scene segmentation.
@@
 The rendered views display the reconstructed bike and guitar captured within the trajectory.
-Our method outperforms baselines by a large margin primarily due to the integration of key-frame optimization and semantic constraints.
+Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints.
 Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
@@
 The bike, garbage bin, and guitar are accurately rendered, meanwhile details are preserved.
-Our method facilitates the selection of key frames base on geometric and semantic constraints, incorporating an uncertainty weighting during the optimization of selected key frames.
-This strategy demonstrate its effectiveness in map optimization from different views meanwhile preventing the unreliable key frame with high uncertainty to significantly altering the earlier accurately reconstructed map.
+Our method facilitates the selection of keyframes based on geometric and semantic constraints, incorporating uncertainty weighting during the optimization of selected keyframes.
+This strategy demonstrates its effectiveness in map optimization from different views meanwhile preventing the unreliable key frame with high uncertainty to significantly altering the earlier accurately reconstructed map.
 4.5 Scene Manipulation
 The obtained semantic mask within the 3D scene has a range of applications for subsequent tasks.
-As an illustrative example, we demonstrate an straightforward but efficient Gaussian editing method defined by eq.
+As an illustrative example, we demonstrate a straightforward but efficient Gaussian editing method defined by eq.
 10, which is crucial for enabling scene manipulation for robotics or mixed reality applications.
@@
 We show the capabilities for object removal and transformation by specifying semantic labels.
-SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a table, as well as moving and rotating the table together with all objects on it.
-Utilizing the decoupled scene representation, in contrast to NeRF-based approaches that demand fine-tuning of the entire network, SGS-SLAM has the ability to choose specific objects within the scene by referencing their semantic mask while keeping the remainder of the well-trained, irrelevant environment fixed.
+SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a jar and flowers, as well as moving and rotating them.
+Utilizing the decoupled scene representation, in contrast to NeRF-based approaches that demand fine-tuning of the entire network, SGS-SLAM can edit specific objects within the scene while keeping the remainder of the well-trained, irrelevant environment fixed.
 As shown in fig.
-4, for object removal, we can directly erase the Gaussians associated with the editing target, such has removing the table while preserving all the items on it.
-In addition, we can group objects by selecting their semantic masks and apply translation and rotation, such as moving and rotating both the table and the above objects to a different place.
+4, we can directly manipulate the Gaussians associated with the editing target, such as erasing, moving, and rotating the jar and flowers on the table.
+In addition, we can group objects by selecting their semantic masks and applying a transition, such as rotating both the table and the above objects as shown in the supplementary material.
 This editing capability requires no training or fine-tuning, making it readily available for downstream applications.
@@
 We perform the ablation of GS-SLAM on the scene0000 of the ScanNet dataset Dai et al.
-(2017a) to evaluate the effectiveness of multi-channel feature supervision, and the key-frame optimization strategies.
+(2017a) to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
 4.6.1 Effect of Multi-channel Optimization
-table 4 shows the ablation study on multi-channel parameter optimization employed in SGS-SLAM.
+table 4 shows the ablation study on multi-channel parameter optimization.
 The results reveal that our optimization strategy can significantly improve the localization and mapping performance.
-Specifically, the system without appearance color, although performs poorly, still maintains its capability for tracking and depth estimation.
-This can be attributed to the gradient flow resulting from semantic supervision.
-Despite the Gaussian rasterizer used in our system not supporting backward flow through depth rendering, this underscores the value of semantic supervision within our framework.
+Specifically, the system without appearance color cannot provide rendered views, whereas camera pose and depth can still be estimated by leveraging depth and semantic input.
 The absence of depth data leads to the poorest depth estimation, highlighting the importance of geometric supervision.
@@
 Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
-w/o color image 7.44 24.59 68.19
-w/o depth map 47.66 40.47 15.14 54.52
-w/o semantic map 9.15 13.81 17.52
-w/o silhouette threshold 29.12 357.48 12.06 28.07
-w/ multi-channel inputs 6.66 11.26 19.47 70.27
-4.6.2 Effect of Key-frame Optimization
-table 5 presents the results of key-frame selection ablation.
-Our two-level key-frame selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance.
-Additionally, without incorporating uncertainty weighting, the system demonstrates decrease in performance compared to its full implementation.
-Table 5: Ablation study of key-frame optimization on the scene0000 of the ScanNet dataset Dai et al.
+without color image 7.44 24.59 68.19
+without depth map 47.66 40.47 15.14 54.52
+without semantic map 9.15 13.81 17.52
+without silhouette threshold 29.12 357.48 12.06 28.07
+with multi-channel inputs 6.66 11.26 19.47 70.27
+4.6.2 Effect of Keyframe Optimization
+table 5 presents the results of keyframe selection ablation.
+Our two-level keyframe selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance.
+Additionally, without incorporating uncertainty weighting, the system demonstrates a decrease in performance compared to its full implementation.
+Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset Dai et al.
 (2017a).
@@
 Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
-w/o geometric threshold 6.66 15.55 19.21 68.93
-w/o semantic threshold 8.44 12.89 17.84 69.85
-w/o uncertainty weighting 6.87 11.43 18.72 70.12
-w/ key-frame selection 6.18 11.26 19.47 70.27
+without geometric threshold 6.66 15.55 19.21 68.93
+without semantic threshold 8.44 12.89 17.84 69.85
+without uncertainty weighting 6.87 11.43 18.72 70.12
+with keyframe selection 6.18 11.26 19.47 70.27
 5 Conclusion and Limitations
 We presented SGS-SLAM, the first semantic dense visual SLAM system based on the 3D Gaussian representation.
-We propose to leverage multi-channel parameter optimization where appearance, geometric, and semantic constraints are combined to enforce high-accurate 3D semantic segmentation, and high-fidelity dense map reconstruction meanwhile effectively produce a robust camera pose estimation.
-SGS-SLAM takes the advantages of optimal key-frame optimization, resulting on reliable reconstruction quality.
-Extensive experiments show that our method provide state-of-the-art tracking and mapping results, meanwhile maintain rapid rendering speeds.
-Furthermore, the high-quality reconstruction of scenes and precise 3D semantic labeling generated by our system establish a strong foundation for the downstream tasks such as scene editing, offers solid prior for the robotics or AR/VR applications.
+We propose to leverage multi-channel parameter optimization where appearance, geometric, and semantic constraints are combined to enforce high-accurate 3D semantic segmentation, and high-fidelity dense map reconstruction meanwhile effectively producing a robust camera pose estimation.
+SGS-SLAM takes advantage of optimal keyframe optimization, resulting in reliable reconstruction quality.
+Extensive experiments show that our method provides state-of-the-art tracking and mapping results, meanwhile maintaining rapid rendering speeds.
+Furthermore, the high-quality reconstruction of scenes and precise 3D semantic labeling generated by our system establish a strong foundation for downstream tasks such as scene editing, offering solid prior for robotics or mixed reality applications.
 5.1 Limitations
@@
 In scenarios where this information is scarce or difficult to access, the system’s effectiveness will be compromised.
-Additionally, our method incurs large memory consumption when deployed in large scenes.
+Additionally, our method incurs large memory consumption when deployed to large scenes.
 Addressing these limitations will be an objective for future research.
 Appendix A Experiment Settings
-In this section, we outline the experimental setup and hyperparameters applied in our studies.
-A.1 SGS-SLAM Settings
+In this section, we outline the experimental setup and hyperparameters applied in our studies. The experiments were conducted on NVIDIA A100-40GB GPUs.
+A.0.1 SGS-SLAM
 By default, both mapping and tracking operations are conducted for each frame.
@@
 The geometric overlap threshold, $\eta$, is defined at 0.05, and the semantic mean Intersection over Union (mIoU) threshold, $T_{\rm sem}$, at 0.7.
-The maximum number of key-frames per frame is limited to 25, considering the computation speed.
+The maximum number of keyframes per frame is limited to 25, considering the computation speed.
 The uncertainty decay coefficient, $\tau$ scales with the length of the input frame series.
@@
 The weights of photometric loss are set to $\lambda_{D}=1.0$, $\lambda_{C}=0.5$, and $\lambda_{S}=0.1$.
-Here, camera parameters are fixed, and Gaussian parameters are optimized, with specific learning rates for 3D position at 1e-4, color 2.5e-3, Gaussian rotation at 1e-3, logit opacity at 0.05, and logg scale at 1e-3.
+Here, camera parameters are fixed, and Gaussian parameters are optimized, with specific learning rates for 3D position at 1e-4, color 2.5e-3, Gaussian rotation at 1e-3, logit opacity at 0.05, and log scale at 1e-3.
 Performance metrics of tracking and mapping are assessed every 5 frames, with mIoU scores evaluated at the same frequency.
@@
 [2023], where the camera transition is large between each frame, the tracking and mapping iterations are adjusted to 220 and 50.
-A.2 Baseline Settings
+A.0.2 Baselines
 We adhere to the default configurations for each baseline as reported in their papers.
@@
 [2023a] are obtained from its paper.
-The frames of training view are chosen based on the results presented in DNS-SLAM.
+The frames of the training view are chosen based on the results presented in DNS-SLAM.
 Compared to NeRF-based models, our approach delivers segmentation results with higher accuracy.
@@
 The ground-truth novel views are captured from meshes.
-SGS-SLAM exhibits rendering of high fidelity, outperforms the NeRF-based ESLAM Johari et al.
+SGS-SLAM exhibits rendering of high fidelity and outperforms the NeRF-based ESLAM Johari et al.
 [2023].
@@
 7.
-Note that the rendering results exhibit minor variations in scene objects due to the use of an modified semantic dataset from ScantNet.
+Note that the rendering results exhibit minor variations in scene objects due to the use of a modified semantic dataset from ScantNet.
 For our method, the training data is processed from the filtered semantic labels using the nyu40-class, where certain objects are not distinctly labeled and are assigned as background (depicted in black).
@@
 In addition, our method also offers accurate instance-level segmentation outcomes.
+B.4 Scene Manipulation
+In this section, we visualize scene manipulation results by grouping the Gaussians using the semantic mask.
+As shown in fig.
+10, for object removal, we can directly erase the Gaussians associated with the editing target, such as removing the table while preserving all the items on it.
+In addition, we can group objects by selecting their semantic masks and applying translation and rotation, such as moving and rotating both the table and the above objects to a different place.
+Figure 10: The visualization of scene manipulation by grouping Gaussians via semantic labels.
+SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a table, as well as moving and rotating the table together with all objects on it.

### v4

@@
-SGS-SLAM: Semantic Gaussian Splatting for Neural Dense SLAM
+
 Abstract
@@
 Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, precise semantic segmentation, and object-level geometric accuracy, while ensuring real-time rendering capabilities.
-Keywords SLAM  $\cdot$ 3D Reconstruction  $\cdot$ 3D Semantic Segmentation
+Keywords:
+SLAM 3D Reconstruction 3D Semantics 3D Segmentation
+**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video can be found at https://youtu.be/y83yw1E-oUo.
+1 Introduction
 Figure 1: The illustration of the proposed SGS-SLAM.
@@
 During the mapping process, SGS-SLAM maps the 2D semantic prior to the 3D scene, jointly optimizing it via the mapping loss for accurate 3D segmentation outcomes.
-**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author
-1 Introduction
 Dense Visual Simultaneous Localization and Mapping (SLAM) is a crucial problem in the field of computer vision.
 It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses in a real-time manner.
-Traditional visual SLAM systems Davison et al.
-(2007); Newcombe et al.
-(2011); Salas-Moreno et al.
-(2013); Mur-Artal et al.
-(2015) stand out in sparse reconstruction using point clouds and voxels, but fall short in dense reconstruction.
-To extract dense geometric information for high-quality representation, learning-based SLAM methods Bloesch et al.
-(2018); Sucar et al.
-(2020) have gained wild attention.
+Traditional visual SLAM systems [6, 29, 32, 27, 39, 9] stand out in sparse reconstruction using point clouds and voxels, but fall short in dense reconstruction.
+To extract dense geometric information for high-quality representation, learning-based SLAM methods [1, 36] have gained wild attention.
 They demonstrate proficiency in generating decent 3D global maps meanwhile exhibiting robustness on noises and outliers.
-In addition, drawing inspiration from the advancements in the neural radiance field (NeRF) Mildenhall et al.
-(2021), NeRF-based SLAM approaches Sucar et al.
-(2021); Zhu et al.
-(2022); Kong et al.
-(2023); Zhang et al.
-(2023); Li et al.
-(2023a); Wang et al.
-(2023) have made further progress.
+In addition, drawing inspiration from the advancements in the neural radiance field (NeRF) [25], NeRF-based SLAM approaches [35, 47, 18, 44, 19, 37, 8] have made further progress.
 They excel in producing accurate and high-fidelity global reconstruction by capturing dense photometric information through differentiable rendering.
@@
 Since the entire scene is modeled through one or several MLPs, it necessitates extensive model tuning for adding or updating scenes.
-In this context, as opposite to NeRF-based neural representation, our exploration shifts towards the volumetric representation based on the 3D Gaussian Radiance Field Kerbl et al.
-(2023).
+In this context, as opposite to NeRF-based neural representation, our exploration shifts towards the volumetric representation based on the 3D Gaussian Radiance Field [17].
 This approach marks a significant shift and offers notable advantages in the scene representation.
@@
 This integration allows real-time rendering of appearance, depth, and semantic color.
-When compared with neural implicit semantic SLAM systems, such as DNS-SLAM Li et al.
-(2023a) and SNI-SLAM Zhu et al.
-(2023), our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy.
+When compared with neural implicit semantic SLAM systems, such as DNS-SLAM [19] and SNI-SLAM [46], our system demonstrates remarkable superiority in terms of rendering speed, reconstruction quality, and segmentation accuracy.
 Leveraging these benefits, our method enables precise editing and manipulation of specific scene elements while preserving the high fidelity of the overall rendering.
@@
 Extensive experiments are conducted on both synthetic and real-world scene benchmarks.
-These experiments compare our method against both implicit NeRF-based approaches Yang et al.
-(2022); Zhu et al.
-(2022); Wang et al.
-(2023); Johari et al.
-(2023), and novel 3D-Gaussian-based methods Keetha et al.
-(2023), evaluating performance in mapping, tracking, and semantic segmentation.
+These experiments compare our method against both implicit NeRF-based approaches [41, 47, 37, 15], and novel 3D-Gaussian-based methods [16], evaluating performance in mapping, tracking, and semantic segmentation.
 Overall, our work presents several key contributions, summarized as follows:
@@
 2.1 Semantic SLAM
-Semantic information is of great importance for SLAM systems Mur-Artal et al.
-(2015); He et al.
-(2023); Whelan et al.
-(2015); Qin et al.
-(2018), which is a crucial requirement for applications in robotics and VR or AR fields.
-Real-time dense semantic SLAM systems Salas-Moreno et al.
-(2013); Bloesch et al.
-(2018); Rosinol et al.
-(2020) integrate semantic information into 3D geometric representations.
-Traditional semantic SLAM systems rely on sparse 3D semantic expressions, such as voxel Hermans et al.
-(2014), point cloud Narita et al.
-(2019), and signed distance field Narita et al.
-(2019).
+Semantic information is of great importance for SLAM systems [27, 12, 38, 30], which is a crucial requirement for applications in robotics and VR or AR fields.
+Real-time dense semantic SLAM systems [32, 1, 31] integrate semantic information into 3D geometric representations.
+Traditional semantic SLAM systems rely on sparse 3D semantic expressions, such as voxel [13], point cloud [28], and signed distance field [28].
 These methods struggle with accurately interpreting complex environments due to limited semantic understanding.
@@
 2.2 Neural Implicit SLAM
-Methods based on NeRF McCormac et al.
-(2018), which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods Chung et al.
-(2023); Li et al.
-(2023b, 2024); Zhou et al.
-(2024).
-iMAP Sucar et al.
-(2021) uses a single MLP for scene representation, which shows limitations in large-scale scenes.
-NICE-SLAM Zhu et al.
-(2022) uses pre-trained multiple MLPs for hierarchical scene representation.
-Co-SLAM Wang et al.
-(2023) combines pixel set-based keyframe tracking with one-blob encoding.
-Go-SLAM Zhang et al.
-(2023) uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding Müller et al.
-(2022) for mapping meanwhile integrated loop-closure detection and global optimization.
+Methods based on NeRF [24], which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods [3, 21, 7, 20, 45].
+iMAP [35] uses a single MLP for scene representation, which shows limitations in large-scale scenes.
+NICE-SLAM [47] uses pre-trained multiple MLPs for hierarchical scene representation.
+Co-SLAM [37] combines pixel set-based keyframe tracking with one-blob encoding.
+Go-SLAM [44] uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding [26] for mapping meanwhile integrated loop-closure detection and global optimization.
 However, these methods cannot utilize semantic information in maps.
-NIDS-SLAM Haghighi et al.
-(2023) leverages the mature front-end tracking of ORB-SLAM3 Campos et al.
-(2021) and Instant-NGP Müller et al.
-(2022) for mapping but does not optimize joint semantic features for 3D reconstruction.
-DNS-SLAM Li et al.
-(2023a) proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features.
-SNI-SLAM Zhu et al.
-(2023), a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
+NIDS-SLAM [11] leverages the mature front-end tracking of ORB-SLAM3 [2] and Instant-NGP [26] for mapping but does not optimize joint semantic features for 3D reconstruction.
+DNS-SLAM [19] proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features.
+SNI-SLAM [46], a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
 2.3 3D Gaussian Splatting SLAM
-The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting Kerbl et al.
-(2023) enable higher efficiency and accuracy on scene reconstruction.
-However, existing 3DGS-based SLAM systems Yan et al.
-(2023); Keetha et al.
-(2023); Huang et al.
-(2023) lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes.
+The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting [17] enable higher efficiency and accuracy on scene reconstruction.
+However, existing 3D-Gaussian-based SLAM systems [40, 16, 14] lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes.
 We fuse semantic features into geometry and appearance and integrate 3D semantic features during keyframe selection in the tracking process.
@@
 SGS-SLAM is a Gaussian-based semantic visual SLAM system.
-section 3.1 introduces its multi-channel Gaussian representation for joint parameter optimization.
+Sec.
+3.1 introduces its multi-channel Gaussian representation for joint parameter optimization.
 Like previous SLAM techniques, our method can be split into two processes: tracking and mapping.
@@
 Mapping optimizes the scene representations based on the estimated camera pose.
-section 3.2 explains the breakdown steps in detail.
-In addition, section 3.3 presents scene manipulation as a case study for downstream tasks.
+Sec.
+3.2 explains the breakdown steps in detail.
+In addition, Sec.
+3.3 presents scene manipulation as a case study for downstream tasks.
 Fig.
@@
 3.1 Multi-Channel Gaussian Representation
-The scene is represented using a Gaussian influence function $f(\cdot)$ on the map, For simplicity, these Gaussians are isotropic, as proposed in Keetha et al. (2023):
+The scene is represented using a Gaussian influence function $f(\cdot)$ on the map, For simplicity, these Gaussians are isotropic, as proposed in [16]:
 $f^{\rm 3D}(x)=\sigma\exp\left(-\frac{\|x-\mu\|^{2}}{2r^{2}}\right)$ (1)
@@
 In order to optimize the parameters of Gaussians to represent the scene, we need to render the Gaussians into 2D images in a differentiable manner.
-We use the render method from Luiten et al.
-(2024), providing extended functionality of rendering depth in colors.
+We use the render method from [22], providing extended functionality of rendering depth in colors.
 It works by splatting 3D Gaussians into the image plane via approximating the integral projection of the influence function $f(\cdot)$ along the depth dimension in pixel coordinates.
@@
 where $K$ is the camera intrinsic matrix, $E_{t}$ is the extrinsic matrix capturing the rotation and translation of the camera at frame $t$, $l$ is the focal length.
-The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula Max (1995):
+The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula [23]:
 $C_{\rm pix}=\sum_{i=1}^{n}c_{i}f_{i,\rm pix}^{\rm 2D}\prod_{j=1}^{i-1}(1-f_{j,\rm pix}^{\rm 2D})$ (3)
@@
 where $d_{i}$ denotes the depth of each Gaussian.
-By setting $d_{i}=1$, we can calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view.
+By setting $d_{i}=1$, we can calculate a silhouette, $Sil_{\rm pix}=D_{\rm pix}(d_{i}=1)$, which assists in determining whether a pixel is visible in the current view [16].
 This aspect of visibility is essential for camera pose estimation, as it relies on the current reconstructed map.
@@
 After densification, the parameters of the map are optimized by minimizing the mapping loss:
-$\mathcal{L}_{\rm mapping}=\mathcal{U}\sum_{\rm pix}\lambda_{D}|D_{\rm pix}^{GT}-D_{\rm pix}|+\lambda_{C}\mathcal{L}_{C}+\lambda_{S}\mathcal{L}_{S}$ (8)
-where $\mathcal{L}_{C}$ and $\mathcal{L}_{S}$ are weighted SSIM loss Kerbl et al. (2023) with respect to appearance image and semantic image:
+$\mathcal{L}_{\rm mapping}=\mathcal{U}_{t}\sum_{\rm pix}\lambda_{D}|D_{\rm pix}^{GT}-D_{\rm pix}|+\lambda_{C}\mathcal{L}_{C}+\lambda_{S}\mathcal{L}_{S}$ (8)
+where $\mathcal{L}_{C}$ and $\mathcal{L}_{S}$ are weighted SSIM loss [17] with respect to appearance image and semantic image:
 $\mathcal{L}(I_{\rm pix})=\sum_{\rm pix}\alpha|I_{\rm pix}^{GT}-I_{\rm pix}|+(1-\alpha)(1-ssim(I_{\rm pix}^{GT},I_{\rm pix}))$ (9)
-$\lambda_{D}$, $\lambda_{C}$, $\lambda_{S}$, and $\alpha$ are predefined hyperparameters, and $\mathcal{U}$ is the uncertainty score defined in section 3.2.2.
-Compared to existing NeRF-based approaches Zhu et al.
-(2022); Johari et al.
-(2023); Li et al.
-(2023a); Zhu et al.
-(2023) that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping.
+$\lambda_{D}$, $\lambda_{C}$, $\lambda_{S}$, and $\alpha$ are predefined hyperparameters, and $\mathcal{U}_{t}$ is the uncertainty score defined in Sec. 3.2.2.
+Compared to existing NeRF-based approaches [47, 15, 19, 46] that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping.
 This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality.
-Compared to recent Gaussian-based methods Keetha et al.
-(2023); Yan et al.
-(2023), SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering.
+Compared to recent Gaussian-based methods [16, 40], SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering.
 This enables the joint optimization of parameters across different channels, remarkably enhancing the efficiency and effectiveness of both mapping and segmentation processes.
@@
 We evaluate our method on both synthetic and real-world datasets.
-To compare with other neural implicit SLAM methods, we evaluate synthetic scenes from Replica dataset Straub et al.
-(2019) and real-world scenes from ScanNet Dai et al.
-(2017a) and ScanNet++ Yeshwanth et al.
-(2023) datasets.
-The ground-truth camera pose and semantic map of Replica are offered from simulation, and the ground-truth camera pose of ScanNet is generated by BundleFusion Dai et al.
-(2017b).
+To compare with other neural implicit SLAM methods, we evaluate synthetic scenes from Replica dataset [34] and real-world scenes from ScanNet [4] and ScanNet++ [43] datasets.
+The ground-truth camera pose and semantic map of Replica are offered from simulation, and the ground-truth camera pose of ScanNet is generated by BundleFusion [5].
 The ground-truth 2D semantic label is provided by the dataset.
@@
 4.1.3 Baselines
-We compare the tracking and mapping with state-of-the-art methods iMap Sucar et al.
-(2021), NICE-SLAM Zhu et al.
-(2022), Co-SLAM Wang et al.
-(2023), ESLAM Johari et al.
-(2023), and SplaTAM Keetha et al.
-(2023).
-For semantic segmentation accuracy, we compare with NIDS-SLAM Haghighi et al.
-(2023), DNS-SLAM Li et al.
-(2023a), and SNI-SLAM Zhu et al.
-(2023).
+We compare the tracking and mapping with state-of-the-art methods iMap [35], NICE-SLAM [47], Co-SLAM [37], ESLAM [15], and SplaTAM [16].
+For semantic segmentation accuracy, we compare with NIDS-SLAM [11], DNS-SLAM [19], and SNI-SLAM [46].
 4.2 Evaluation of Mapping and Localization
-We show quantitative measures of reconstruction quality using the Replica dataset Straub et al.
-(2019) in table 1.
+We show quantitative measures of reconstruction quality using the Replica dataset [34] in Tab.
+1.
 Our method demonstrates state-of-the-art performance.
 Compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
-Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset Straub et al.
-(2019), with key details accentuated using colorful boxes.
+Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset [34], with key details accentuated using colorful boxes.
 The results demonstrate that our method delivers more high-fidelity and robust reconstructions, particularly by capturing more detailed features of the objects in the scene.
@@
 In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes.
-Compared with SplaTAM, which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes.
+Compared with SplaTAM [16], which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes.
 This combination enables our method to achieve thorough modeling of both objects and their surrounding environment.
 The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
-Table 1: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset Straub et al. (2019). Our method demonstrates SOTA performances on all three metrics.
-Methods Metrics Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
+Table 1: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset [34]. Our method demonstrates SOTA performances in most cases among three metrics.
+  Methods Metrics Avg.
+Room0
+Room1
+Room2
+Office0
+Office1
+Office2
+Office3
+Office4
+[…]
+24.42
+22.12
+22.47
+24.52
+29.07
+30.34
+19.66
+22.23
+24.94
+NICE-SLAM
+[…]
+0.809
+0.689
+0.757
+0.814
+0.874
+0.886
+0.797
+0.801
+0.856
+[…]
+0.233
+0.330
+0.271
+0.208
+0.229
+0.181
+0.235
+0.209
+0.198
+[…]
+30.24
+27.27
+28.45
+29.06
+34.14
+34.87
+28.43
+28.76
+30.91
+Co-SLAM
+[…]
+0.939
+0.910
+0.909
+0.932
+0.961
+0.969
+0.938
+0.941
+0.955
+[…]
+0.252
+0.324
+0.294
+0.266
+0.209
+0.196
+0.258
+0.229
+0.236
+[…]
+29.08
+25.32
+27.77
+29.08
+33.71
+30.20
+28.09
+28.77
+29.71
+ESLAM
+[…]
+0.929
+0.875
+0.902
+0.932
+0.960
+0.923
+0.943
+0.948
+0.945
+[…]
+0.336
+0.313
+0.298
+0.248
+0.184
+0.228
+0.241
+0.196
+0.204
+[…]
+33.98
+32.48
+33.72
+34.96
+38.34
+39.04
+31.90
+29.70
+31.68
+SplaTAM
+[…]
+0.969
+0.975
+0.970
+0.982
+0.982
+0.982
+0.965
+0.950
+0.946
+[…]
+0.099
+0.072
+0.096
+0.074
+0.083
+0.093
+0.100
+0.118
+0.155
+[…]
+34.66 32.50 34.25 35.10 38.54 39.20 32.90 32.05 32.75
+Ours
+[…]
+0.973 0.976 0.978
+0.981
+0.984
+0.980
+0.967 0.966 0.949
+[…]
+0.096 0.070 0.094 0.070
+0.086
+0.087
+0.101
+0.115 0.148
+Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset [34].
+The values represent the average outcomes across eight scenes.
+The results of baselines are retrieved from [46].
+Our method remarkably outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics.
+Note that for parameter size, we only count the parameters of the model.
+SOTA performances are highlighted.
+  Methods
+Depth L1 [cm]$\downarrow$
+ATE Mean [cm]$\downarrow$
+ATE RMSE [cm]$\downarrow$
 […]
@@
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset Straub et al.
-(2019).
-The values represent the average outcomes across eight scenes.
-The results of baselines are retrieved from Zhu et al.
-(2023).
-Our method remarkably outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics.
-Note that for parameter size, we only count the parameters of the model.
-[…]
-iMAP 4.645 3.118 4.153 9.92 2.23 1.82
-[…]
-[…]
-[…]
-[…]
-table 2 displays the tracking evaluation results on the Replica dataset Straub et al.
-(2019).
+iMAP
+4.645
+3.118
+4.153
+9.92
+2.23
+1.82
+NICE-SLAM
+1.903
+1.795
+2.503
+13.70
+0.20
+0.20
+Co-SLAM
+1.513
+0.935
+1.059
+17.24
+10.20 6.41
+ESLAM
+1.180
+0.520
+0.630
+18.11
+3.62
+3.02
+Ours 0.356 0.327 0.412
+5.27
+3.52
+2.11
+Tab.
+2 displays the tracking evaluation results on the Replica dataset [34].
 Our method excels in achieving the highest level of depth L1 loss (cm) and minimal ATE error, surpassing baseline methods by 70% in terms of depth loss and 34% in terms of ATE RMSE (cm).
@@
 4.3 Evaluation of Semantic Segmentation
-table 3 shows a quantitative evaluation of our method in comparison to other neural semantic SLAM approaches.
+Tab.
+3 shows a quantitative evaluation of our method in comparison to other neural semantic SLAM approaches.
 It’s worth noting that we only show four scenes because previous NeRF-based semantic models only reported results on these scenes.
@@
 Our approach offers an unparalleled ability to identify 3D objects in decomposed representations, which can serve as 3D priors for tracking and mapping in future time steps, and is well-suited for further downstream tasks.
-Table 3: Quantitative comparison of our method against existing semantic NeRF-based SLAM methods on the Replica dataset Straub et al.
-(2019).
+Table 3: Quantitative comparison of our method against existing semantic NeRF-based SLAM methods on the Replica dataset [34].
 The baselines are limited to four scenes as their results are reported only for these.
 For each scene, we compute the average mIoU score by comparing the rendered and the ground-truth 2D semantic image in the training view.
-Our method significantly outperforms the NeRF-based approaches, achieving mIoU scores over 90%.
-Methods Avg. mIoU [%$\uparrow$ Room0 [%$\uparrow$ Room1 [%$\uparrow$ Room2 [%$\uparrow$ Office0 [%$\uparrow$
-NIDS-SLAM 82.37 82.45 84.08 76.99 85.94
-[…]
-[…]
-[…]
+Our method significantly outperforms the NeRF-based approaches, achieving SOTA mIoU scores over 90%.
+  Methods Avg. mIoUnormal-$\uparrow$
+Room0 [%$\uparrow$
+Room1 [%$\uparrow$
+Room2 [%$\uparrow$
+Office0 [%$\uparrow$
+NIDS-SLAM
+82.37
+82.45
+84.08
+76.99
+85.94
+DNS-SLAM
+84.77
+88.32
+84.90
+81.20
+84.66
+SNI-SLAM
+87.41
+88.42
+87.43
+86.16
+87.63
+Ours 92.72 92.95 92.91 92.10 92.90
 4.4 Evaluation of Keyframe Optimization
-Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset Dai et al.
-(2017a).
-The rendered views display the reconstructed bike and guitar captured within the trajectory.
+Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset [4].
+The rendered views display the reconstructed objects such as bike, fridge, garbage bin, and guitar from novel views.
 Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints.
@@
 In real-world datasets, tracking errors tend to accumulate along a trajectory, making pose estimations at later timestamps less reliable.
-Such inaccuracies can compromise the quality of map reconstructions, negatively impacting the accuracy of a previously well-established scene.
-A case in point is scene0000 from the ScanNet dataset Dai et al.
-(2017a), where objects such as bike and guitar are revisited at early and late stages in the trajectory.
-Keyframes from later in the sequence, influenced by inaccurate camera poses, can disrupt the previously accurate reconstructions of these objects.
+Such inaccuracies can compromise the quality of map reconstructions, negatively impacting the previously well-established scenes.
+A case in point is scene0000 from the ScanNet dataset [4], where objects such as bike and guitar are revisited at early and late stages in the trajectory.
+Keyframes from later sequences, influenced by inaccurate camera poses, can disrupt the previously accurate reconstructions.
 Fig.
-3 illustrates the novel-view evaluation for scene0000.
-In comparison to SplaTAM, depicted in the upper row, our method delivers more accurate reconstruction outcomes.
+3 illustrates the novel view evaluation for scene0000.
+In comparison to ESLAM [15] and SplaTAM [16], which are based on NeRF and 3D Gaussians, our method delivers more accurate reconstruction outcomes.
 The bike, garbage bin, and guitar are accurately rendered, meanwhile details are preserved.
 Our method facilitates the selection of keyframes based on geometric and semantic constraints, incorporating uncertainty weighting during the optimization of selected keyframes.
-This strategy demonstrates its effectiveness in map optimization from different views meanwhile preventing the unreliable key frame with high uncertainty to significantly altering the earlier accurately reconstructed map.
+This strategy demonstrates its effectiveness in map optimization from different views meanwhile preventing the unreliable keyframse with high uncertainty to significantly altering the earlier accurately reconstructed map.
 4.5 Scene Manipulation
@@
 10, which is crucial for enabling scene manipulation for robotics or mixed reality applications.
-Figure 4: The case study on scene manipulation in room0 of the Replica dataset Straub et al.
-(2019).
+Figure 4: The case study on scene manipulation in room0 of the Replica dataset [34].
 We show the capabilities for object removal and transformation by specifying semantic labels.
@@
 4.6 Ablation Study
-We perform the ablation of GS-SLAM on the scene0000 of the ScanNet dataset Dai et al.
-(2017a) to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
+We perform the ablation of SGS-SLAM on the scene0000 of the ScanNet dataset [4] to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
 4.6.1 Effect of Multi-channel Optimization
-table 4 shows the ablation study on multi-channel parameter optimization.
+Tab.
+4 shows the ablation study on multi-channel parameter optimization.
 The results reveal that our optimization strategy can significantly improve the localization and mapping performance.
@@
 Without this threshold, the system shows a significant decline in the effectiveness of tracking and mapping.
-Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset Dai et al.
-(2017a).
+Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset [4].
 The comparison involves settings where appearance, depth, and semantic supervision are removed.
+means the metric is inapplicable.
   Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
-without color image 7.44 24.59 68.19
-without depth map 47.66 40.47 15.14 54.52
-without semantic map 9.15 13.81 17.52
-without silhouette threshold 29.12 357.48 12.06 28.07
-with multi-channel inputs 6.66 11.26 19.47 70.27
+without color image ($C_{\rm pix}$) 7.44 24.59 68.19
+without depth map ($D_{\rm pix}$) 47.66 40.47 15.14 54.52
+without semantic map ($S_{\rm pix}$) 9.15 13.81 17.52
+without silhouette threshold ($Sil_{\rm pix}$) 29.12 357.48 12.06 28.07
+with multi-channel optimization 6.18 11.26 19.47 70.27
 4.6.2 Effect of Keyframe Optimization
-table 5 presents the results of keyframe selection ablation.
+Tab.
+5 presents the results of keyframe selection ablation.
 Our two-level keyframe selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance.
 Additionally, without incorporating uncertainty weighting, the system demonstrates a decrease in performance compared to its full implementation.
-Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset Dai et al.
-(2017a).
-The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
+Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset [4]. The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
   Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
-without geometric threshold 6.66 15.55 19.21 68.93
-without semantic threshold 8.44 12.89 17.84 69.85
-without uncertainty weighting 6.87 11.43 18.72 70.12
+without geometric threshold ($T_{\rm geo}$) 6.66 15.55 19.21 68.93
+without semantic threshold ($T_{\rm sem}$) 8.44 12.89 17.84 69.85
+without uncertainty weighting ($\mathcal{U}$) 6.87 11.43 18.72 70.12
 with keyframe selection 6.18 11.26 19.47 70.27
@@
 Furthermore, the high-quality reconstruction of scenes and precise 3D semantic labeling generated by our system establish a strong foundation for downstream tasks such as scene editing, offering solid prior for robotics or mixed reality applications.
-5.1 Limitations
+5.0.1 Limitations
 SGS-SLAM replies on depth and 2D semantic signal inputs for tracking and mapping.
@@
 Addressing these limitations will be an objective for future research.
-Appendix A Experiment Settings
-In this section, we outline the experimental setup and hyperparameters applied in our studies. The experiments were conducted on NVIDIA A100-40GB GPUs.
-A.0.1 SGS-SLAM
-By default, both mapping and tracking operations are conducted for each frame.
-During the tracking phase, we set the silhouette visibility threshold, $T_{\rm sil}$, to 0.99.
-The multi-channel optimization involves three parameters: $\lambda_{D}=1.0$ for depth, $\lambda_{C}=0.5$ for colors, and $\lambda_{S}=0.05$ for semantic loss, with the semantic loss weight being comparatively low due to the typical noisiness of real-world semantic labels.
-Throughout the tracking, the multi-channel Gaussian parameters remain constant, adjusting only the camera parameters with a learning rate of 2e-3 for transition.
-Key-frames are initially chosen at intervals of every 5 frames, then refined based on geometric and semantic criteria.
-The geometric overlap threshold, $\eta$, is defined at 0.05, and the semantic mean Intersection over Union (mIoU) threshold, $T_{\rm sem}$, at 0.7.
-The maximum number of keyframes per frame is limited to 25, considering the computation speed.
-The uncertainty decay coefficient, $\tau$ scales with the length of the input frame series.
-In mapping process, the silhouette threshold $T_{\rm sil}$ is adjusted to 0.5.
-The weights of photometric loss are set to $\lambda_{D}=1.0$, $\lambda_{C}=0.5$, and $\lambda_{S}=0.1$.
-Here, camera parameters are fixed, and Gaussian parameters are optimized, with specific learning rates for 3D position at 1e-4, color 2.5e-3, Gaussian rotation at 1e-3, logit opacity at 0.05, and log scale at 1e-3.
-Performance metrics of tracking and mapping are assessed every 5 frames, with mIoU scores evaluated at the same frequency.
-The mapping and tracking iteration steps are specific to each dataset, In the case of the Replica dataset Straub et al.
-[2019], the number of iterations for tracking and mapping are set to 40 and 60.
-For the ScanNet dataset Dai et al.
-[2017a], tracking and mapping are set to 120 and 40.
-In the enhanced ScanNet++ dataset Yeshwanth et al.
-[2023], where the camera transition is large between each frame, the tracking and mapping iterations are adjusted to 220 and 50.
-A.0.2 Baselines
-We adhere to the default configurations for each baseline as reported in their papers.
-The evaluation metrics for tracking and mapping are consistent with those applied to our method.
-For baselines whose implementations are not publicly available, we present the results as reported in their papers.
-Appendix B Additional Experiment Results
-We provide additional quantitative analysis of camera tracking in section B.1.
-The visualization of semantic segmentation compared with NeRF-based method is presented in section B.2.
-More qualitative novel view rendering results are illustrated in section B.3.
-We compared our method with Vox-Fusion Yang et al.
-[2022], NICE-SLAM Zhu et al.
-[2022], Co-SLAM Wang et al.
-[2023], ESLAM Johari et al.
-[2023], and Point-SLAM Sandström et al.
-[2023] for ATE RMSE evaluation.
-For 3D semantic segmentation, we visualized the comparison with DNS-SLAM Li et al.
-[2023a].
-B.1 Camera Tracking
-In this section, we break down the quantitative analysis on ATE RMSE [cm] on Replica Straub et al.
-[2019], ScanNet Dai et al.
-[2017a], and ScanNet++ Yeshwanth et al.
-[2023] datasets.
-table 6, table 7, and table 8 present the evaluation our SGS-SLAM against baseline models on each dataset.
-Our method of estimating camera poses by directly optimizing the gradient on dense photometric loss achieves state-of-the-art tracking performance on datasets with high-quality RGB-D images.
-In particular, on the ScanNet++ dataset Yeshwanth et al.
-[2023], where there is a large camera transition between successive frames, NeRF-based methods like ESLAM failed to track.
-Conversely, SGS-SLAM demonstrated robust tracking performance.
-Table 6: Quantitative comparison of ATE RMSE [cm] between our method and the baselines for each scene of the Replica dataset Straub et al. [2019].
-Methods Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
-Vox-Fusion 3.09 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 7: Quantitative comparison of ATE RMSE [cm] between our method and the baselines for the selected scenes on the ScanNet dataset Dai et al. [2017a].
-Methods Avg. $0$ 0059 0106 0169 0181 0207
-Vox-Fusion 26.90 68.84 24.18 8.41 27.28 23.30 9.41
-[…]
-[…]
-[…]
-[…]
-Table 8: Quantitative comparison of ATE RMSE [cm] between our method and the baseline for the selected scenes on the ScanNet++ dataset Yeshwanth et al. [2023].
-Methods Avg. [cm]$\downarrow$ 8b5caf3398 [cm]$\downarrow$ b20a261fdf [cm]$\downarrow$
-ESLAM 170.06 185.15 156.96
-[…]
-B.2 Semantic Segmentation
-In this section, the outcomes of semantic segmentation on the Replica dataset Straub et al.
-[2019] are visualized and compared with DNS-SLAM Li et al.
-[2023a], a NeRF-based approach.
-As illustrated, our method offered accurate and detailed segmentation, whereas DNS-SLAM faces challenges in edges due to the over-smoothing issue of NeRF.
-Figure 5: Qualitative comparison of our method and DNS-SLAM Li et al.
-[2023a] for semantic segmentation from the Replica dataset Straub et al.
-[2019].
-The visualization outcomes of DNS-SLAM Li et al.
-[2023a] are obtained from its paper.
-The frames of the training view are chosen based on the results presented in DNS-SLAM.
-Compared to NeRF-based models, our approach delivers segmentation results with higher accuracy.
-B.3 Novel View Rendering
-We present additional results of novel view rendering using our method across the Replica Straub et al.
-[2019], ScanNet Dai et al.
-[2017a], and ScanNet++ Yeshwanth et al.
-[2023] datasets, with comparisons to ESLAM Johari et al.
-[2023].
-Visualizations are provided in fig.
-6, fig.
-7, fig.
-8, and fig.
-9 with semantic segmentation outcomes.
-Our method consistently delivers high-quality rendering results for both synthesized and real-world datasets.
-Notably, on the challenging real-world ScanNet++ dataset, ESLAM Johari et al.
-[2023] struggled to reconstruct the scene.
-By contrast, SGS-SLAM provides accurate high-fidelity scene reconstructions along with precise segmentation outcomes.
-Note that the ground-truth segmentation labels are retrieved from the ground-truth mesh at the instance level, and therefore, our results also show instance-level segmentation.
-Figure 6: The visualization of novel view rendering between the ESLAM Johari et al.
-[2023] and our method on the Replica dataset Straub et al.
-[2019].
-The ground-truth novel views are captured from meshes provided by the dataset.
-Figure 7: The visualization of novel view rendering between the baseline and our method using the ScanNet dataset Dai et al.
-[2017a].
-The ground-truth novel views are captured from meshes.
-SGS-SLAM exhibits rendering of high fidelity and outperforms the NeRF-based ESLAM Johari et al.
-[2023].
-In contrast to the ground-truth mesh, our method demonstrates robust mapping in areas where the ground-truth mesh presents holes.
-Figure 8: The visualization of 3D semantic segmentation results of SGS-SLAM, as applied to the novel views selected in fig.
-7.
-Note that the rendering results exhibit minor variations in scene objects due to the use of a modified semantic dataset from ScantNet.
-For our method, the training data is processed from the filtered semantic labels using the nyu40-class, where certain objects are not distinctly labeled and are assigned as background (depicted in black).
-Furthermore, we introduce extra labels, like guitar, bag, and basket, to enhance the quality of scene reconstruction.
-Figure 9: The visualization of novel view rendering between the baseline and our method using the ScanNet++ dataset Yeshwanth et al.
-[2023].
-The ground-truth novel views are captured from meshes.
-SGS-SLAM demonstrates superior rendering quality, while ESLAM Johari et al.
-[2023] suffers from significant tracking errors and fails to reconstruct the map.
-In addition, our method also offers accurate instance-level segmentation outcomes.
-B.4 Scene Manipulation
-In this section, we visualize scene manipulation results by grouping the Gaussians using the semantic mask.
-As shown in fig.
-10, for object removal, we can directly erase the Gaussians associated with the editing target, such as removing the table while preserving all the items on it.
-In addition, we can group objects by selecting their semantic masks and applying translation and rotation, such as moving and rotating both the table and the above objects to a different place.
-Figure 10: The visualization of scene manipulation by grouping Gaussians via semantic labels.
-SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a table, as well as moving and rotating the table together with all objects on it.

### v5

@@
 SLAM 3D Reconstruction 3D Semantics 3D Segmentation
-**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video can be found at https://youtu.be/y83yw1E-oUo.
+**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video is available at https://youtu.be/y83yw1E-oUo.
 1 Introduction
@@
 Table 1: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset [34]. Our method demonstrates SOTA performances in most cases among three metrics.
-  Methods Metrics Avg.
-Room0
-Room1
-Room2
-Office0
-Office1
-Office2
-Office3
-Office4
+  Methods Metrics Avg. Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4
 […]
-24.42
-22.12
-22.47
-24.52
-29.07
-30.34
-19.66
-22.23
-24.94
-NICE-SLAM
 […]
-0.809
-0.689
-0.757
-0.814
-0.874
-0.886
-0.797
-0.801
-0.856
 […]
-0.233
-0.330
-0.271
-0.208
-0.229
-0.181
-0.235
-0.209
-0.198
 […]
-30.24
-27.27
-28.45
-29.06
-34.14
-34.87
-28.43
-28.76
-30.91
-Co-SLAM
 […]
-0.939
-0.910
-0.909
-0.932
-0.961
-0.969
-0.938
-0.941
-0.955
 […]
-0.252
-0.324
-0.294
-0.266
-0.209
-0.196
-0.258
-0.229
-0.236
 […]
-29.08
-25.32
-27.77
-29.08
-33.71
-30.20
-28.09
-28.77
-29.71
-ESLAM
 […]
-0.929
-0.875
-0.902
-0.932
-0.960
-0.923
-0.943
-0.948
-0.945
 […]
-0.336
-0.313
-0.298
-0.248
-0.184
-0.228
-0.241
-0.196
-0.204
 […]
-33.98
-32.48
-33.72
-34.96
-38.34
-39.04
-31.90
-29.70
-31.68
-SplaTAM
 […]
-0.969
-0.975
-0.970
-0.982
-0.982
-0.982
-0.965
-0.950
-0.946
 […]
-0.099
-0.072
-0.096
-0.074
-0.083
-0.093
-0.100
-0.118
-0.155
 […]
-34.66 32.50 34.25 35.10 38.54 39.20 32.90 32.05 32.75
-Ours
 […]
-0.973 0.976 0.978
-0.981
-0.984
-0.980
-0.967 0.966 0.949
 […]
-0.096 0.070 0.094 0.070
-0.086
-0.087
-0.101
-0.115 0.148
 Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset [34].
@@
 SOTA performances are highlighted.
-  Methods
-Depth L1 [cm]$\downarrow$
-ATE Mean [cm]$\downarrow$
-ATE RMSE [cm]$\downarrow$
 […]
-[…]
-[…]
-iMAP
-4.645
-3.118
-4.153
-9.92
-2.23
-1.82
-NICE-SLAM
-1.903
-1.795
-2.503
-13.70
-0.20
-0.20
-Co-SLAM
-1.513
-0.935
-1.059
-17.24
-10.20 6.41
-ESLAM
-1.180
-0.520
-0.630
-18.11
-3.62
-3.02
-Ours 0.356 0.327 0.412
-5.27
-3.52
-2.11
+iMAP 4.645 3.118 4.153 9.92 2.23 1.82
+NICE-SLAM 1.903 1.795 2.503 13.70 0.20 0.20
+Co-SLAM 1.513 0.935 1.059 17.24 10.20 6.41
+ESLAM 1.180 0.520 0.630 18.11 3.62 3.02
+Ours 0.356 0.327 0.412 5.27 3.52 2.11
 Tab.
@@
 Our method significantly outperforms the NeRF-based approaches, achieving SOTA mIoU scores over 90%.
-  Methods Avg. mIoUnormal-$\uparrow$
-Room0 [%$\uparrow$
-Room1 [%$\uparrow$
-Room2 [%$\uparrow$
-Office0 [%$\uparrow$
-NIDS-SLAM
-82.37
-82.45
-84.08
-76.99
-85.94
-DNS-SLAM
-84.77
-88.32
-84.90
-81.20
-84.66
-SNI-SLAM
-87.41
-88.42
-87.43
-86.16
-87.63
+[…]
+NIDS-SLAM 82.37 82.45 84.08 76.99 85.94
+DNS-SLAM 84.77 88.32 84.90 81.20 84.66
+SNI-SLAM 87.41 88.42 87.43 86.16 87.63
 Ours 92.72 92.95 92.91 92.10 92.90
@@
 We perform the ablation of SGS-SLAM on the scene0000 of the ScanNet dataset [4] to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
+Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset [4].
+The comparison involves settings where appearance, depth, and semantic supervision are removed.
+means the metric is inapplicable.
+  Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
+without color image ($C_{\rm pix}$) 7.44 24.59 68.19
+without depth map ($D_{\rm pix}$) 47.66 40.47 15.14 54.52
+without semantic map ($S_{\rm pix}$) 9.15 13.81 17.52
+without silhouette threshold ($Sil_{\rm pix}$) 29.12 357.48 12.06 28.07
+with multi-channel optimization 6.18 11.26 19.47 70.27
+Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset [4]. The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
+  Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
+without geometric threshold ($T_{\rm geo}$) 6.66 15.55 19.21 68.93
+without semantic threshold ($T_{\rm sem}$) 8.44 12.89 17.84 69.85
+without uncertainty weighting ($\mathcal{U}$) 6.87 11.43 18.72 70.12
+with keyframe selection 6.18 11.26 19.47 70.27
 4.6.1 Effect of Multi-channel Optimization
@@
 Without this threshold, the system shows a significant decline in the effectiveness of tracking and mapping.
-Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset [4].
-The comparison involves settings where appearance, depth, and semantic supervision are removed.
-means the metric is inapplicable.
-  Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
-without color image ($C_{\rm pix}$) 7.44 24.59 68.19
-without depth map ($D_{\rm pix}$) 47.66 40.47 15.14 54.52
-without semantic map ($S_{\rm pix}$) 9.15 13.81 17.52
-without silhouette threshold ($Sil_{\rm pix}$) 29.12 357.48 12.06 28.07
-with multi-channel optimization 6.18 11.26 19.47 70.27
 4.6.2 Effect of Keyframe Optimization
@@
 Additionally, without incorporating uncertainty weighting, the system demonstrates a decrease in performance compared to its full implementation.
-Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset [4]. The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
-  Settings Depth L1 [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ PSNR [dB]$\uparrow$ mIoU [%$\uparrow$
-without geometric threshold ($T_{\rm geo}$) 6.66 15.55 19.21 68.93
-without semantic threshold ($T_{\rm sem}$) 8.44 12.89 17.84 69.85
-without uncertainty weighting ($\mathcal{U}$) 6.87 11.43 18.72 70.12
-with keyframe selection 6.18 11.26 19.47 70.27
 5 Conclusion and Limitations

### v6

@@
-
+SGS-SLAM: Semantic Gaussian Splatting For Neural Dense SLAM
 Abstract
@@
 Extensive experiments demonstrate that SGS-SLAM delivers state-of-the-art performance in camera pose estimation, map reconstruction, precise semantic segmentation, and object-level geometric accuracy, while ensuring real-time rendering capabilities.
+The implementation code is available at https://github.com/ShuhongLL/SGS-SLAM.
 Keywords:
-SLAM 3D Reconstruction 3D Semantics 3D Segmentation
-**footnotetext: These authors contributed equally to this workfootnotetext: Corresponding author  footnotetext: The demo video is available at https://youtu.be/y83yw1E-oUo.
+SLAM 3D Reconstruction 3D Segmentation
+  footnotetext: The demo video is available at https://youtu.be/y83yw1E-oUo.
 1 Introduction
@@
 Dense Visual Simultaneous Localization and Mapping (SLAM) is a crucial problem in the field of computer vision.
-It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses in a real-time manner.
-Traditional visual SLAM systems [6, 29, 32, 27, 39, 9] stand out in sparse reconstruction using point clouds and voxels, but fall short in dense reconstruction.
+It aims to reconstruct a dense 3D map in an unseen environment while simultaneously tracking the camera poses.
+Traditional visual SLAM systems [6, 32, 35, 30] stand out in sparse mapping using point clouds and voxels, but fall short in dense reconstruction.
 To extract dense geometric information for high-quality representation, learning-based SLAM methods [1, 39] have gained wild attention.
 They demonstrate proficiency in generating decent 3D global maps meanwhile exhibiting robustness on noises and outliers.
-In addition, drawing inspiration from the advancements in the neural radiance field (NeRF) [25], NeRF-based SLAM approaches [35, 47, 18, 44, 19, 37, 8] have made further progress.
+Drawing inspiration from advancements in the neural radiance field (NeRF) [28], NeRF-based SLAM approaches [38, 51, 20, 48, 21, 41, 8] have made further progress.
 They excel in producing accurate and high-fidelity global reconstruction by capturing dense photometric information through differentiable rendering.
@@
 Benefits from its rasterization of 3D primitives, Gaussian Splatting exhibits remarkably fast rendering speeds and allows direct gradient flow to each Gaussian’s parameters.
-This results in an almost linear projection between the dense photometric loss and parameters during optimization, unlike the hierarchical pixel sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models.
-Moreover, the direct projection capability simplifies the addition of new parameters to the Gaussian field as separate channels, thereby enabling dynamic multi-channel rendering.
+This results in an almost linear projection between the dense photometric loss and parameters during optimization [18], unlike the hierarchical pixel sampling and indirect gradient flow through multiple non-linear layers seen in NeRF models.
+Moreover, the direct projection capability simplifies the addition of new channels to the Gaussian field, thereby enabling dynamic multi-channel feature rendering.
 Crucially, we integrate a semantic map into the 3D Gaussian field, essential for applications in robotics and mixed reality.
@@
 Overall, our work presents several key contributions, summarized as follows:
-  • We introduce SGS-SLAM, the first semantic dense visual SLAM system grounded in 3D Gaussians.
+  • We introduce SGS-SLAM, the first semantic RGB-D SLAM system grounded in 3D Gaussians.
 SGS-SLAM employs an explicit volumetric representation, enabling swift and real-time camera tracking and scene mapping.
@@
 Furthermore, SGS-SLAM utilizes these diverse channels for keyframe selection during the tracking phase, concentrating on actively recognizing objects seen earlier in the trajectory.
-This approach results in efficient and high-quality map reconstruction based on the chosen keyframes.
-  • Utilizing semantic representation, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation.
-SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects in the map in real time.
-This is achieved by grouping Gaussians by specifying the semantic labels of the objects.
+  • Utilizing the semantic map, SGS-SLAM provides a highly accurate disentangled object representation in 3D scenes, laying a solid foundation for downstream tasks such as scene editing and manipulation.
+SGS-SLAM facilitates the dynamic moving, rotating, or removal of objects that is achieved by grouping Gaussians by specifying the semantic labels of the objects.
 2 Related Work
@@
 2.2 Neural Implicit SLAM
-Methods based on NeRF [24], which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods [3, 21, 7, 20, 45].
+Methods based on NeRF [27], which handle complex topological structures and differentiable scene representation methods, have garnered significant attention, leading to the development of neural implicit SLAM methods [3, 11, 23, 22, 9, 49, 10].
 iMAP [38] uses a single MLP for scene representation, which shows limitations in large-scale scenes.
 NICE-SLAM [51] uses pre-trained multiple MLPs for hierarchical scene representation.
-Co-SLAM [37] combines pixel set-based keyframe tracking with one-blob encoding.
-Go-SLAM [44] uses Droid-SLAM as a front-end tracking system and multi-resolution hash encoding [26] for mapping meanwhile integrated loop-closure detection and global optimization.
-However, these methods cannot utilize semantic information in maps.
-NIDS-SLAM [11] leverages the mature front-end tracking of ORB-SLAM3 [2] and Instant-NGP [26] for mapping but does not optimize joint semantic features for 3D reconstruction.
+Co-SLAM [41] combines pixel-set-based keyframe tracking with one-blob encoding.
+Go-SLAM [48] uses Droid-SLAM [40] as the tracking system and multi-resolution hash encoding [29] for mapping.
+However, these methods cannot utilize semantic information in the map.
+NIDS-SLAM [13] leverages the tracking system of ORB-SLAM3 [2] and Instant-NGP [29] for mapping but does not optimize joint semantic features for 3D reconstruction.
 DNS-SLAM [21] proposes a 2D semantic prior system that provides multi-view geometry constraints but does not optimize 3D reconstruction with semantic features.
-SNI-SLAM [46], a work parallel to ours, introduces semantic loss for geometric supervision but remains limited by the efficiency constraints of NeRF’s volume rendering.
+DNS-SLAM [21] and SNI-SLAM [50] introduce semantic loss for geometric supervision but remain limited by the efficiency constraints of NeRF’s volume rendering.
 2.3 3D Gaussian Splatting SLAM
-The outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting [17] enable higher efficiency and accuracy on scene reconstruction.
-However, existing 3D-Gaussian-based SLAM systems [40, 16, 14] lack traditional effective tracking optimization, limiting tracking accuracy and the ability to recognize semantic information in scenes.
-We fuse semantic features into geometry and appearance and integrate 3D semantic features during keyframe selection in the tracking process.
-This allows us to obtain more effective and higher-resolution scene segmentation results while maintaining real-time performance.
+Utilizing the outstanding performance and fast rasterization capabilities of 3D Gaussian Splatting [19], Gaussian-based SLAM systems offer higher efficiency and accuracy on scene reconstruction [43, 18, 16, 26, 47, 7, 24].
+However, existing Gaussian-based SLAM systems lack the ability to recognize semantic information in scenes.
+To bridge this gap, We utilize the semantic map during keyframe selection and integrate the semantic feature loss in the tracking and mapping process.
+This allows us to obtain more effective and higher-quality scene segmentation outcomes meanwhile preserving real-time processing performance.
 3 Method
@@
 where $K$ is the camera intrinsic matrix, $E_{t}$ is the extrinsic matrix capturing the rotation and translation of the camera at frame $t$, $l$ is the focal length.
-The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering using the Max volume rendering formula [23]:
+The influence of all Gaussians on this pixel can be combined by sorting the Gaussians in depth order and performing front-to-back volume rendering:
 $C_{\rm pix}=\sum_{i=1}^{n}c_{i}f_{i,\rm pix}^{\rm 2D}\prod_{j=1}^{i-1}(1-f_{j,\rm pix}^{\rm 2D})$ (3)
@@
 Specifically, we randomly select pixels from the current frame and extract their corresponding Gaussians $G_{\rm sample}$ in the 3D scene.
-These Gaussians, $G_{\rm sample}$, are then projected onto the camera views of keyframes as $G_{\rm proj}$.
-The $G_{\rm proj}$ are evaluated based on the geometric overlap ratio:
+These sampled $G_{\rm sample}$ are then projected onto the camera views of keyframes as $G_{\rm proj}$, which are evaluated based on the geometric overlap ratio:
 $\eta=\frac{1}{\sum G_{\rm proj}}\sum_{n=i}\{G_{i}|0\leq width(G_{i})\leq W,0\leq height(G_{i})\leq H\}$ (7)
@@
 $\lambda_{D}$, $\lambda_{C}$, $\lambda_{S}$, and $\alpha$ are predefined hyperparameters, and $\mathcal{U}_{t}$ is the uncertainty score defined in Sec. 3.2.2.
-Compared to existing NeRF-based approaches [47, 15, 19, 46] that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping.
-This approach allows for direct gradient flow to each parameter, resulting in high rendering speeds and optimal reconstruction quality.
+Compared to existing NeRF-based approaches [51, 17, 21, 50] that necessitate complex model architectures and feature fusion strategies, SGS-SLAM adopts explicit Gaussian representation for mapping, resulting in high rendering speeds and optimal reconstruction quality.
 Compared to recent Gaussian-based methods [18, 43], SGS-SLAM incorporates geometric, appearance, and semantic features for multi-channel rendering.
@@
 4.1.2 Metrics
-We use PSNR, Depth-L1 (on 2D depth maps), SSIM, and LPIPS to evaluate the reconstruction quality.
-For the evaluation of camera pose, we adopt the average absolute trajectory error (ATE RMSE).
+We use PSNR, Depth-L1, SSIM, and LPIPS to evaluate the reconstruction quality.
+To evaluate the camera pose, we use the average absolute trajectory error (ATE RMSE).
 For semantic segmentation, we calculate mIoU score.
 4.1.3 Baselines
-We compare the tracking and mapping with state-of-the-art methods iMap [35], NICE-SLAM [47], Co-SLAM [37], ESLAM [15], and SplaTAM [16].
+We compare the tracking and mapping with state-of-the-art methods NICE-SLAM [51], vMap [20], Co-SLAM [41], ESLAM [17], and SplaTAM [18].
 For semantic segmentation accuracy, we compare with NIDS-SLAM [13], DNS-SLAM [21], and SNI-SLAM [50].
+Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset [37], with key details accentuated using colored boxes.
+The results demonstrate that our method delivers more high-fidelity and robust reconstructions.
 4.2 Evaluation of Mapping and Localization
@@
 Compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
-Figure 2: Qualitative comparison of our method and the baselines for reconstruction across three scenes from the Replica Dataset [34], with key details accentuated using colorful boxes.
-The results demonstrate that our method delivers more high-fidelity and robust reconstructions, particularly by capturing more detailed features of the objects in the scene.
-In Fig.
-2, we present the reconstruction results of three chosen scenes, where regions of interest are accentuated with boxes in various colors.
-Our method exhibits high-fidelity reconstruction outcomes.
-Specifically, for small, intricately textured objects like a clock, socket, books on a tea table, and a lamp, our approach shows remarkable accuracy over NeRF-based methods.
-This is because Gaussians are capable of representing objects with complex textures and surfaces.
-Furthermore, NeRF-based methods often struggle with the over-smoothing issue, resulting in blurred edges on objects.
-In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes.
-Compared with SplaTAM [16], which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes.
-This combination enables our method to achieve thorough modeling of both objects and their surrounding environment.
-The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
 Table 1: Quantitative comparison of our method and the baselines in training view rendering on the Replica dataset [37]. Our method demonstrates SOTA performances in most cases among three metrics.
@@
 […]
-Table 2: Quantitative comparison in terms of Depth L1, ATE, memory usage, and FPS between our method and the baselines on the Replica dataset [34].
+Table 2: Quantitative comparison in terms of Depth L1, ATE, and FPS between our method and the baselines on the Replica dataset [37].
 The values represent the average outcomes across eight scenes.
+The FPS is evaluated by setting the same number of training iterations for all systems for fair comparison.
 The results of baselines are retrieved from [50].
-Our method remarkably outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics.
-Note that for parameter size, we only count the parameters of the model.
+Our method outperforms the baselines at Depth and ATE evaluations, and performs fairly on FPS metrics.
 SOTA performances are highlighted.
-[…]
-iMAP 4.645 3.118 4.153 9.92 2.23 1.82
+  Methods Depth L1 [cm]$\downarrow$ ATE Mean [cm]$\downarrow$ ATE RMSE [cm]$\downarrow$ Track. FPS [f/s]$\uparrow$ Map. FPS [f/s]$\uparrow$ SLAM FPS [f/s]$\uparrow$
 NICE-SLAM 1.903 1.795 2.503 13.70 0.20 0.20
@@
 ESLAM 1.180 0.520 0.630 18.11 3.62 3.02
+SplaTAM 0.525 0.348 0.454 5.53 3.84 2.26
 Ours 0.356 0.327 0.412 5.27 3.52 2.11
+In Fig.
+2, we present the reconstruction results of three chosen scenes, where regions of interest are accentuated with boxes in various colors.
+Our method exhibits high-fidelity reconstruction outcomes.
+Specifically, for small, intricately textured objects like a clock, socket, books on a tea table, and a lamp, our approach shows remarkable accuracy over NeRF-based methods.
+This is because Gaussians are capable of representing objects with complex textures and surfaces.
+Furthermore, NeRF-based methods often struggle with the over-smoothing issue, resulting in blurred edges on objects.
+In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes.
+Compared with SplaTAM [18], which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes.
+This combination enables our method to achieve thorough modeling of both objects and their surrounding environment.
+The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
 Tab.
@@
 4.4 Evaluation of Keyframe Optimization
-Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset [4].
+Figure 3: The selected novel view synthesis of scene0000 from the ScanNet dataset [4].
 The rendered views display the reconstructed objects such as bike, fridge, garbage bin, and guitar from novel views.
 Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints.
-Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
+Note that the ground-truth for novel views is captured from the offline-reconstructed mesh provided by the ScanNet dataset.
 In real-world datasets, tracking errors tend to accumulate along a trajectory, making pose estimations at later timestamps less reliable.
@@
 4.6 Ablation Study
-We perform the ablation of SGS-SLAM on the scene0000 of the ScanNet dataset [4] to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
-Table 4: Ablation study of multi-channel optimization on the scene0000 of the ScanNet dataset [4].
+We perform the ablation of SGS-SLAM on the scene0000_00 of the ScanNet dataset [4] to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
+Table 4: Ablation study of multi-channel optimization on the scene0000_00 of the ScanNet dataset [4].
 The comparison involves settings where appearance, depth, and semantic supervision are removed.
@@
 with multi-channel optimization 6.18 11.26 19.47 70.27
+4.6.1 Effect of Multi-channel Optimization
 Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset [4]. The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
@@
 with keyframe selection 6.18 11.26 19.47 70.27
-4.6.1 Effect of Multi-channel Optimization
 Tab.